### v1

# TempoControl: Temporal Attention Guidance for Text-to-Video Models

Abstract
Recent advances in generative video models have enabled the creation of high-quality videos based on natural language prompts. However, these models frequently lack fine-grained temporal control, meaning they do not allow users to specify when particular visual elements should appear within a generated sequence. In this work, we introduce TempoControl, a method that allows for temporal alignment of visual concepts during inference, without requiring retraining or additional supervision. TempoControl utilizes cross-attention maps, a key component of text-to-video diffusion models, to guide the timing of concepts through a novel optimization approach. Our method steers attention using three complementary principles: aligning its temporal shape with a control signal (via correlation), amplifying it where visibility is required (via energy), and maintaining spatial focus (via entropy). TempoControl allows precise control over timing while ensuring high video quality and diversity. We demonstrate its effectiveness across various video generation applications, including temporal reordering for single and multiple objects, as well as action and audio-aligned generation.
Figure 1: Applications of our inference-time temporal control method (TempoControl), which enables multi-object and single-object control, motion control, and audio alignment. Words that are controlled to appear throughout the entire video are shown in blue. Words that appear according to a temporally varying control signal are shown in orange.
Codehttps://github.com/Shira-Schiber/TempoControl
Project Pagehttps://shira-schiber.github.io/TempoControl/
1 Introduction
Generative video models have improved dramatically, enabling the synthesis of high-quality, temporally coherent videos from natural language descriptions (Singer et al. 2022; Blattmann et al. 2023; Yang et al. 2024; Bar-Tal et al. 2024; Wan et al. 2025). These models have enabled creative applications in animation, design, and virtual content creation by generating visually compelling sequences that align with textual descriptions. However, while current models offer strong spatial consistency and global semantics, they lack fine-grained temporal control. Specifically, guiding the model to generate visual elements at specific time points is particularly challenging. This includes making an object appear in the middle of a scene or aligning visual aspects of lighting with the sound of thunder (see Figure 1).
Recent works on generative video modeling have proposed various spatial and motion control mechanisms, including object masking, layout conditioning, or added camera controls (Gupta et al. 2024; Wang et al. 2024b; Feng et al. 2025; He et al. 2024). In contrast, temporal control has received significantly less attention. Incorporating such control typically requires additional supervision in the form of temporally annotated datasets. Since video data is often scarce, extending annotations to capture temporal variation is expensive and often impractical at scale. Synthetic data could be used; however, generating diverse and temporally accurate video-text pairs, especially for abstract concepts such as movement, is a highly non-trivial challenge.
In this work, we propose a lightweight yet effective inference-time control method, TempoControl, which introduces temporal control for generated video. Instead of retraining the generative model, TempoControl leverages a property already embedded in these systems: the text-frame cross-attention mechanism.
Cross-attention layers inherently encode strong signals about which specific words in the prompt are realized in each generated frame. By directly influencing these attention maps during the denoising process, we can shift the temporal alignment of visual concepts without sacrificing quality or requiring additional data. The approach is data-efficient and preserves both the fidelity and diversity of the original generation process.
Attention control is achieved by carefully steering the latent variables during the diffusion process. At each denoising step, we apply a few stochastic gradient descent iterations (Battash, Wolf, and Lindenbaum 2024) until a satisfactory level of temporal alignment is achieved, notably, without updating model parameters.
To guide temporal alignment, TempoControl introduces a novel loss with three complementary components, based on: Pearson correlation, energy, and entropy. The Pearson term encourages the temporal shape of attention to match that of an external control signal, operating on the normalized attention vector to align the presence of a concept with the desired timing. However, since correlation is scale-invariant, it may yield high scores even when attention values are too low to render the object visible. To mitigate this, we incorporate an energy term that directly promotes stronger attention in frames where the temporal signal is high, and suppresses it elsewhere. Additionally, we introduce an entropy regularization term to maintain spatial focus, ensuring that when attention is activated, it remains coherent and not diffusely spread across the frame.
Beyond single-object scenarios, as shown in Figure 1, our method generalizes well to multi-object scenes. In such cases, it is important to maintain the visual grounding of non-target entities; we find that preserving one object (e.g., the bird) while selectively manipulating another (e.g., the cat) helps retain the overall scene structure. In addition to nouns, our method can also manipulate verbs (e.g., clapping), effectively modifying the motion depicted in the video. Finally, we explore the potential of aligning external audio cues with video generation. While we focus on simple cases involving a single audio event (e.g., a lightning strike), we observe promising results, suggesting that our approach may naturally extend to zero-shot multimodal control without additional training.
Our contributions are as follows: (i) We propose an inference-time method for temporally controlling the appearance of specific words in text-to-video diffusion models, without requiring model finetuning or additional supervision. (ii) We introduce a suite of losses that operate over spatiotemporal attention to enable fine-grained control over both the timing and intensity of word grounding. (iii) We demonstrate the versatility of our method across various settings, including multi-object control, motion manipulation, and audio-to-video alignment, and introduce a new metric to evaluate the task of temporal alignment control, which is not covered by existing benchmark suites.
Figure 2: Video generated for the prompt “A cat and a dog.” Before (left) and after optimization (right). Top: Spatial attention maps for the tokens cat and dog at denoising step t=3t{=}3, shown for frames j=2,10,19j{=}2,10,19. Attention is overlaid on generated frames. Bottom: temporal attention $a_{i,j}^{t}$ (blue) vs. target mask $m_{i,j}$ (orange), with Pearson correlation loss shown.
2 Related Work
Motion Control. Fine-grained motion control has been an active area of research. Boximator (Wang et al. 2024b) introduces control by allowing users to specify the spatial locations of a start box and an end box for each object. This control is implemented by fine-tuning a dedicated control module. BlobGEN-Vid (Feng et al. 2025) enhances control by introducing visual primitive-based guidance, such as lines, to enable detailed manipulation of object motion, fine-grained appearance, and smooth object transitions. CameraCtrl (He et al. 2024) addresses controllability by enabling precise camera movement and perspective control. This is accomplished by training a plug-and-play camera pose control module on top of a video diffusion model. Motion Prompting (Geng et al. 2025) enables movement control by incorporating spatio-temporally sparse or dense motion trajectory inputs. However, these approaches require extensive datasets and computational resources for training, substantial user annotations during inference, and notably lack precise control over motion timing or peak movement appearance.
Attention-based Control. Attention-based methods have improved spatial control in text-to-image generation. Prompt-to-Prompt (Hertz et al. 2023) enables word-level edits via cross-attention manipulation, while Attend-and-Excite (Chefer et al. 2023) ensures all mentioned objects are depicted by balancing token attention. Battash et al. (2024) introduced a layout optimization strategy for multi-subject generation in text-to-image diffusion, which constrains cross-attention to improve the spatial separation of objects. Be Yourself (Dahary et al. 2024) uses bounded attention to preserve subject identity and prevent semantic leakage in multi-subject text-to-image generation. Stable Flow (Avrahami et al. 2025) enables training-free image editing by selectively injecting attention features into vital layers of a diffusion transformer. These methods operate on static images, whereas we extend attention-based control to the temporal domain, enabling control over object timing and motion in video generation.
Inference time optimization. Inference-time optimization for text-to-video diffusion has only recently attracted attention. (Li et al. 2025) pre-compute a spatio-temporal sketch and inject plan-consistent noise during sampling to tighten prompt alignment. (Wei et al. 2025) converts a single keyframe into a textured, deformable 3D mesh that is re-rendered at every denoising step to maintain consistency in the garment’s appearance. (Shaulov et al. 2025) derive variance-based optical-flow cues from the model’s latent predictions and add a flow-coherence loss on-the-fly, smoothing motion without retraining. (Lian et al. 2023) let a large language model generate scene graphs and temporal layouts that act as attention masks, improving complex spatio-temporal grounding. While these methods improve fidelity and consistency without necessitating additional training, none offer explicit, user-defined control over the timing of textual element unfolding.
3 Preliminaries
We build TempoControl on top of Wan 2.1 (Wan et al. 2025), a state-of-the-art diffusion model that serves as our backbone for text-to-video generation. Let $v\in\mathbb{R}^{T_{v}\times H_{v}\times W_{v}\times 3}$ represent the input video, where $T_{v}$ is the number of frames, $H_{v}\times W_{v}$ denotes the spatial resolution of each frame, and the final dimension corresponds to the three RGB color channels. The video is first encoded by a 3D causal variational autoencoder into a latent tensor
$z\in\mathbb{R}^{T^{\prime}\times H^{\prime}\times W^{\prime}\times C},$ (1)
where $T^{\prime}$, $H^{\prime}$, and $W^{\prime}$ denote the temporal and spatial dimensions of the latent representation, which are reduced relative to the input video. Flattening the spatial and temporal dimensions, $z$ is reshaped into a sequence of $n_{v}$ video tokens, where $n_{v}=T^{\prime}\cdot H^{\prime}\cdot W^{\prime}$, and each token is a $C$-dimensional feature vector representing a specific spatiotemporal region of the video.
The text prompt $\mathcal{P}=[p_{1},\dots,p_{n_{p}}]$ is embedded using a pretrained text encoder. For brevity, we refer to tokens as words, although a word may consist of multiple subtokens. All operations described for words are applied to the average of their corresponding subtokens. Video synthesis is then performed as a denoising diffusion process in the latent space, where the model iteratively refines latent variables $\{z_{t}\}_{t=T}^{0}$, gradually denoising $z_{T}$ into $z_{0}$. At each timestep $t$, a transformer-based network $\operatorname{DiT}(z_{t},\mathcal{P},t)$ predicts the noise in $z_{t}$, conditioned on the prompt $\mathcal{P}$ and the timestep $t$.
3.1 Cross-attention in text-to-video diffusion models
A key challenge in text-to-video generation is ensuring that the synthesized video aligns semantically and temporally with the input prompt. In diffusion-based architectures, this alignment is achieved through cross-attention layers within the transformer blocks at each denoising step $t$.
During cross-attention, video tokens act as queries that attend to text tokens, which serve as keys and values. This mechanism enables the model to inject contextually relevant linguistic information into each video token representation. The attention weights are computed via scaled dot products between queries and keys, producing a set of attention tensors across layers, $A^{t}\in\mathbb{R}^{L\times h\times n_{v}\times n_{p}}$, where $L$ is the number of cross-attention layers, $h$ is the number of attention heads, $n_{v}$ is the number of video tokens, and $n_{p}$ is the number of text tokens. A softmax is applied along the text-token dimension, yielding a probability distribution over prompt words for each video token, per head and per layer.
To obtain a global view of token-level influence, we average attention maps across heads and cross-attention layers, producing an aggregate attention matrix $\bar{A}^{t}\in\mathbb{R}^{n_{v}\times n_{p}}$. This matrix quantifies the influence of each text token on each video token.
The effect of a specific word $i$ is measured by extracting the corresponding column $\bar{A}^{t}_{i}\in\mathbb{R}^{n_{v}}$. The vector $\bar{A}^{t}_{i}$ can be reshaped at temporal index $j$, yielding $\bar{A}^{t}_{j,i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$, which represents the spatial influence of word $i$ at temporal position $j$ and denoising step $t$. Figure 2 illustrates this for the words cat and dog in the prompt “A cat and a dog.” The attention maps are up-sampled to the frame resolution for visualization.
The attention map captures the spatiotemporal grounding of words in the generated video; TempoControl is specifically designed to steer the temporal influence of a token. Therefore, we focus on the scalar attention value $\hat{A}^{t}_{j,i}=\langle\bar{A}^{t}_{j,i}\rangle_{x,y}\in\mathbb{R}$, where $\langle\cdot\rangle_{x,y}$ denotes summation over the spatial coordinates $x\in[1,H^{\prime}]$ and $y\in[1,W^{\prime}]$. Intuitively, $\hat{A}^{t}_{j,i}$ measures the influence of word $i$ on the latent representation at temporal position $j$ during the denoising step $t$.
To summarize the temporal attention pattern of word $i$ at denoising step $t$, we collect the scalar values across frames into a vector:
$a_{i}^{t}=[\hat{A}^{t}_{1,i},\hat{A}^{t}_{2,i},\ldots,\hat{A}^{t}_{T^{\prime},i}]\in\mathbb{R}^{T^{\prime}}.$ (2)
We omit the first frame due to unstable attention values that do not reliably reflect word grounding. This vector serves as the target for our steering method, described in the next section.
Figure 3: Illustration of TempoControl. During a single denoising step $t$, we extract spatial attention maps $\bar{A}^{t}_{j,i}$ (for word $i$ at temporal index $j$), aggregate to a temporal attention signal $a_{i}^{t}$, and align it with the target mask vector $m_{i}$ via temporal and spatial losses. Gradients $\nabla\mathcal{L}$ are used to update the latent code $z_{t}$.
4 Method
Given a video $v$, our goal is to generate an edited version $v^{\prime}$ such that specific tokens appear according to a temporal condition defined by a weighting mask $m_{i}=[m_{i,1},\ldots,m_{i,T^{\prime}}]$, where $m_{i,j}\in[0,1]$ indicates the desired strength of activation of token $p_{i}$ at temporal position $j$ (see Figure 2 bottom). In most experiments, we use binary masks (i.e., values of zero or one), except for the audio-alignment setting, where the mask reflects sound strength. The main steps of our approach are detailed in the appendix Algorithm 1, illustrated in Figure 3, and discussed below.
We steer the latent representation during the first $k$ denoising steps by performing $l$ gradient update iterations at each step using the AdamW optimizer. For clarity, the update can be expressed as:
$z^{\prime}_{t}=z_{t}-\alpha\nabla_{z_{t}}\mathcal{L}^{t},$ (3)
where $\alpha\in\mathbb{R}$ is the learning rate and $t$ is the current denoising step. The loss $\mathcal{L}^{t}$ consists of four components (detailed below): (i) temporal correlation, (ii) energy, (iii) entropy regularization, and (iv) spatial consistency.
Temporal correlation term.
The first loss term encourages the temporal attention vector $a_{i}^{t}$ to match the target mask $m_{i}$ using the Pearson correlation:
$\mathcal{L}_{\text{corr}}^{t}=-\frac{\text{Cov}(m_{i},\tilde{a}_{i}^{t})}{\sigma_{m_{i}}\sigma_{\tilde{a}_{i}^{t}}},$ (4)
where $\tilde{a}_{i}^{t}$ is the min-max normalized version of $a_{i}^{t}$, defined as $\tilde{a}_{i}^{t}=\frac{a_{i}^{t}-\min(a_{i}^{t})}{\max(a_{i}^{t})-\min(a_{i}^{t})}$. Here, $\text{Cov}(\cdot,\cdot)$ denotes empirical covariance, and $\sigma_{m_{i}}$ and $\sigma_{\tilde{a}_{i}^{t}}$ are the corresponding standard deviations. Note that if $m_{i}$ has zero variance, such as when encouraging an object to appear throughout the entire video, the correlation term is undefined; in such cases, it is excluded from the loss.
This alignment term adjusts the temporal placement of words in the video but does not affect their overall strength. As a result, words that are initially invisible may remain so. The following loss term is designed to modulate the temporal strength of attention.
Attention energy term.
The energy term adjusts the overall magnitude of attention for word token $i$ based on the temporal mask $m_{i}$. We use the indicator function $\mathbb{1}_{\{\cdot\}}$, which equals 1 when the condition is true and 0 otherwise. The positive energy encourages attention where the word should be active:
$\mathcal{L}_{\oplus}^{t}=\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}_{\{m_{i,j}>\tau\}}\cdot a_{i,j}^{t}.$ (5)
The negative energy penalizes attention where the word should not be visible:
$\mathcal{L}_{\ominus}^{t}=\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}_{\{m_{i,j}\leq\tau\}}\cdot a_{i,j}^{t}.$ (6)
The final energy loss is defined as the net difference:
$\mathcal{L}_{\text{energy}}^{t}=\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}.$ (7)
Here, $\tau\in[0,1]$ is a threshold that determines whether a given time index is considered active for token $p_{i}$. While $m_{i}$ is often binary (i.e., 0 or 1), it may also take continuous values (e.g., for soft audio-driven conditioning). In such cases, the indicator $\mathbb{1}_{\{m_{i,j}>\tau\}}$ selects the time-steps where the target activation is sufficiently strong.
Increasing the attention magnitude without constraint can lead to overly diffuse spatial maps (Qi et al. 2025), especially in multi-object scenes, which degrades the visual grounding of word tokens. To regularize this behavior and preserve spatial consistency, we introduce additional loss terms in the following paragraph.
Entropy regularization.
Word-level attention is typically spatially focused rather than dispersed (see Figure 2). To discourage excessive spreading of attention during optimization, we regularize the entropy of the spatial attention maps. Specifically, we apply this regularization only at temporal indices where the word is expected to be active (i.e., $m_{i,j}>\tau$), resulting in:
$\mathcal{L}_{\text{entropy}}^{t}=\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}>\tau]\cdot\mathcal{H}(\bar{A}^{t}_{j,i}).$ (8)
Here, $\mathcal{H}(\bar{A})$ denotes the Shannon entropy computed over the spatial dimensions.
Spatial consistency penalty.
To maintain spatial consistency, we penalize deviations between the initial and optimized spatial attention maps for the target word. Here, $\hat{A}^{t}_{j,i}$ denotes the spatial attention map before optimization at step $t$, for word $i$ at frame $j$. The penalty is defined as
$\mathcal{L}_{\text{spatial}}^{t}=\left\|\sum_{j=1}^{T^{\prime}}\left(\hat{A}^{t}_{j,i}-\bar{A}^{t}_{j,i}\right)\right\|_{2}^{2},$ (9)
where the difference is computed over the spatial dimensions, after summing across all latent frames.
Full objective.
The combined loss at step $t$ is:
$\mathcal{L}^{t}=\mathcal{L}_{\text{corr}}^{t}+\lambda_{1}\mathcal{L}_{\text{energy}}^{t}+\lambda_{2}\mathcal{L}_{\text{entropy}}^{t}+\lambda_{3}\mathcal{L}_{\text{spatial}}^{t},$ (10)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are hyperparameters used to calibrate the contribution of the regularization terms. This objective is optimized over latent codes during inference; model weights remain fixed.
Correlation-based early stopping.
Some videos naturally align more closely with the temporal objective and therefore require fewer optimization steps. To adaptively allocate compute, we introduce a stopping rule based on the Pearson correlation in the temporal loss $\mathcal{L}_{\text{corr}}^{t}$. Specifically, for each denoising step $t$, we perform up to $l$ optimization iterations. If the correlation exceeds a threshold $\tau_{\text{corr}}$, we terminate optimization early for that step; otherwise, we continue refining until the limit is reached.
Method Temporal Accuracy T. Absence Accuracy T. Presence Accuracy Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality One Object Text 63.94% 67.38% 60.50% 84.66% 92.30% 99.16% 47.50% 45.98% 53.76% Ours 81.00% 88.75% 72.91% 80.60% 90.33% 99.05% 38.75% 44.37% 52.99% Two Objects Text 37.50% 45.85% 29.15% 92.06% 95.32% 98.92% 62.20% 53.26% 68.56% Ours 55.00% 66.46% 43.54% 92.21% 95.43% 99.19% 41.46% 56.55% 70.66% Movement Text 19% 93.17% 95.91% 98.10% 73.00% 55.81% 60.46% Ours 53% 92.86% 95.67% 98.20% 80.00% 55.06% 61.76%
Table 1: Comparison against Wan 2.1 across three temporal-control applications: (i) one object, (ii) two objects, and (iii) movement. Results for irrelevant metrics are marked with –. Higher values indicate better performance for all metrics.
Method Multiple Object (GriT) Multiple Object (YOLO) Dynamic Degree Text 74.13% 61.54% 30.49% Ours 76.37% 65.73% 18.78%
Table 2: Multiple object benchmark assessing the ability to depict multiple objects across the video.
Question Type Ours Baseline Same Temporal Accuracy 61.51% 16.94% 21.55% Visual Quality 62.66% 25.99% 11.35%
Table 3: User study results for temporal accuracy and visual quality.
5 Experiments
Below, we describe the applications and evaluation protocols used in our experiments.
Setup.
We use Wan 2.1 1.3B as our text-to-video backbone. All experiments were run on a single NVIDIA H200 GPU. TempoControl takes approximately $\sim$460s per video, compared to $\sim$170s for vanilla generation. Hyperparameter details and ablations are available in the appendix.
Baseline.
We employ Wan 2.1 1.3B with explicit temporal cues in the prompt (e.g., “…at the third second…”). We describe the exact phrasing details in the appendix.
Metrics.
We evaluate on VBench (Huang et al. 2024), reporting temporal-related metrics: Subject Consistency (main subject coherence across frames), Background Consistency (background stability), Motion Smoothness (absence of temporal flicker), and Dynamic Degree (overall motion magnitude), as well as quality metrics: Aesthetic Quality (visual appeal) and Imaging Quality (clarity of frames). Additional details about these metrics are provided in the appendix.
We also introduce a new metric, Temporal Accuracy, which measures whether objects appear in the correct frames as specified by the prompt, and provides breakdowns for both object absence and presence. By default, the official VBench metric uses the GRiT detector (Wu et al. 2024). To provide a more up-to-date assessment, we report temporal accuracy results using YOLOv10 (Wang et al. 2024a).
For movement, Temporal Accuracy is assessed by matching the movement time specified in the prompt with the second exhibiting the most movement, as determined by optical flow. Each video is sampled at 4 frames per second; we compute the average flow magnitude per second, and the second with the highest average is considered the most active.
5.1 Quantitative Results
Single object temporal control.
In Table 1 (One Object), our method improves temporal accuracy by 17.06% to 81.00%, with similar gains in temporal absence (+21.37) and presence (+12.41) accuracy compared to text prompts with explicit time cues. Image Quality and Aesthetic Quality remain comparable, with a difference of 1%.
We observe decreases in Dynamic Degree, dropping by 8.75% to 38.75%, reflecting reduced dynamics due to temporal control, which requires object presence over extended periods of the video.
Two objects temporal control.
In Table 1 (Two Objects), we show that the two-object setting presents a significant challenge, with the baseline achieving only 37.5% temporal accuracy. Our method improves to 55.0%. Temporal absence and presence accuracies also show gains of 20.6% and 14.4%, respectively. Imaging and Aesthetic Quality increase by about 2% each. Subject Consistency, Background Consistency, and Motion Smoothness remain similar between methods. At the same time, we find that Dynamic Degree is reduced here, as in the one-object case, by roughly 21%. This reflects the trade-off imposed by temporal control, which requires static background objects and reduces dynamic content.
Figure 4: Examples of videos generated by Wan 2.1 (left) and our method with temporal conditioning (right) for prompts that specify temporal conditions. The prompts are simplified for brevity. We show frames corresponding to the first, middle, and last third of each video.
Figure 5: Examples of video alignment to an audio signal.
Movement temporal control.
Table 1 (Movement) shows that TempoControl significantly improves temporal accuracy (53% vs. 19%) compared to the text-prompt baseline. We note that the original temporal accuracy of the model is similar to random, indicating that Wan 2.1 does not capture movement cues at all. The Dynamic Degree here increases, as the goal is to introduce more movement, demonstrating that our approach successfully creates more pronounced motion. Imaging and Aesthetic Quality remain on par with the baseline. Overall, our explicit temporal guidance sharpens movement timing without compromising visual quality and consistency.
Multiple objects benchmark.
We also report results on the VBench Multiple Objects benchmark in Table 2, which tests the ability to generate two objects described in the input text throughout the entire video. We observe improved performance on the Multiple Object metric (76.34% vs. 74.13%), although this comes at the cost of reduced video dynamics. Detailed results are provided in the appendix.
Human evaluation.
We conducted a human evaluation with 50 annotators, each comparing 16 videos (8 one-object, 8 two-object) generated by TempoControl and Wan 2.1. Annotators selected which video was better in terms of temporal accuracy and visual quality, or if they were the same. As shown in Table 3, our method outperformed the baseline in temporal accuracy (61.51% vs. 16.94%) and also improved visual quality. For temporal accuracy, 21.55% of responses indicated no preference, likely due to subtle timing differences. Full details of the evaluation are provided in the appendix.
5.2 Qualitative Results
In Figure 4, we show four videos generated to depict objects or actions at specific times. In the first row, where a single apple is expected to appear at the fourth second, Wan 2.1 generates the apple near the middle of the five-second video, while our method correctly places it near the end of the video. For the “cat and dog, dog at first half” prompt, Wan 2.1 fails to show the dog at the beginning of the video, whereas our approach accurately places the dog in the first half. In the “magician pulling a rabbit last second” scenario, Wan 2.1 displays the rabbit from the start, instead of only in the final second as required. Similarly, for “tiger leaping first second,” Wan 2.1 does not capture the leaping motion at the correct time, while our method, with temporal conditioning, accurately generates the requested action. All videos are included in the supplementary material.
5.3 Audio-Visual Alignment
We also experimented with zero-shot alignment of video to external audio by using a preprocessed audio envelope as the temporal condition. Given a waveform $y$, we extract its onset strength envelope $s_{t}$, adaptively pool it to match the number of frames, and normalize to $[0,1]$ via min-max normalization. Prominent peaks are preserved via thresholding and Gaussian smoothing:
$\tilde{s}_{t}=\begin{cases}\hat{s}_{t},&\text{if }\hat{s}_{t}\geq\tau_{audio},\\ \text{GaussianSmooth}(\hat{s}_{t};\sigma),&\text{otherwise}.\end{cases}$ (11)
where $\tau_{audio}$ is a threshold for peak preservation, and $\sigma$ is the standard deviation of the Gaussian kernel used for smoothing.
In Figure 5, we present examples of accurate video adaptation to audio peaks, with additional results in the appendix. Although experiments are limited to simple cases with prominent peaks, these findings suggest a promising direction for aligning generated video with audio without requiring costly training on curated datasets.
5.4 Discussions
Text temporal control limitations.
In our experiments, adding explicit timing to Wan 2.1 prompts did not improve temporal accuracy and often reduced video and image quality. For example, temporal accuracy fell from 65% to 64%, and video quality from 56% to 53%. For fairness, all reported results use the same prompt setup as the text baseline, despite omitting explicit timing, which yields higher performance for our approach. For instance, for a single object, we observe a video quality of 56% compared to 53%, with the same temporal accuracy of 81%. See the appendix for details.
Consistency limitations.
Temporal conditioning may lead to perceptual changes in background or unspecified features. For example, it can alter the color of a cat (see Figure 1). Although perceptual consistency is often preserved, our current objectives do not fully guarantee it, which remains a promising area for future work.
5.5 Conclusions
Current text-to-video models generate semantically accurate and high-quality videos, but they lack precise control over the temporal appearance of objects and actions. Our results show that existing text-conditioning approaches are insufficient for temporal control. To address this gap, we present a data-free temporal conditioning technique, which maintains video quality while enabling fine-grained temporal control through attention steering. We hope our approach will inspire new solutions to tasks such as audio-video alignment, as discussed in this work.
Supplementary Material: TempoControl – Temporal Attention Guidance for Text-to-Video Models
This supplementary material provides detailed information supporting our main paper. It includes algorithmic implementation, experimental design, application-specific configurations, extended quantitative and qualitative results, and a human evaluation setup. Videos and code are provided as part of the supplementary materials package.
Specifically, we include:
  1. 1.
    Full algorithm and pseudocode (Section .6).
  2. 2.
    Experimental setup and hyperparameters (Section .7).
  3. 3.
    Configuration details for each application setting (Section .8).
  4. 4.
    Human evaluation protocol and interface (Section .9).
  5. 5.
    Extended quantitative results for multi-object grounding (Section .10).
  6. 6.
    Ablation study of loss components and qualitative failure modes (Section .11).
  7. 7.
    Analysis of prompt phrasing and its impact on performance (Section .12).
  8. 8.
    Additional qualitative results with sampled frames (Section .13).
.6 Algorithm Pseudo-code
Algorithm 1 outlines our inference-time optimization procedure. Starting from noise $z_{T}$, we iteratively steer the latent code during the first $k$ denoising steps of the diffusion process. At each step $t$, we compute cross-attention maps from the video-text model $\operatorname{DiT}$ and derive the scalar temporal attention $a_{i}^{t}$ for each token $p_{i}$.
The optimization loss $\mathcal{L}$ combines four components: correlation alignment $\mathcal{L}_{\text{c}}^{t}$, energy modulation $\mathcal{L}_{\text{enr}}^{t}$, entropy regularization $\mathcal{L}_{\text{ent}}^{t}$, and spatial consistency $\mathcal{L}_{\text{s}}^{t}$. The latent code is updated via gradient descent for up to $l$ steps per denoising timestep, or until the Pearson correlation exceeds a threshold $\tau_{\text{corr}}$. After optimization, the standard denoising process resumes to produce the final latent $z_{0}$.
Algorithm 1 TempoControl: Temporal Attention-Guided Inference-Time Steering
Input: Prompt $\mathcal{P}$, tokens $p_{i}$ with temporal masks $m_{i}$, total diffusion steps $T$, optimized steps $k$, max updates per step $l$, correlation threshold $\tau_{\text{corr}}$, model $\operatorname{DiT}$
Output: Final latent $z_{0}$
1:Initialize $z_{T}\sim\mathcal{N}(0,I)$
2:for $t=T,T{-}1,\dots,T{-}k{+}1$ do
3:  for $u=1$ to $l$ do
4:   Extract cross-attentions $A^{t}$ from $\operatorname{DiT}(z_{t},\mathcal{P},t)$
5:   Average heads and layers: $\bar{A}^{t}$
6:   $\mathcal{L}\leftarrow 0$
7:   for each $p_{i}$ with $m_{i}$ do
8:     $a_{i}^{t}\leftarrow[\langle\bar{A}^{t}_{j,i}\rangle_{x,y}]_{j=1}^{T^{\prime}}$
9:     $\mathcal{L}_{\text{c}}^{t}\leftarrow-\text{Corr}(m_{i},\text{minmax}(a_{i}^{t}))$
10:     $\mathcal{L}_{\oplus}^{t}\leftarrow\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}>\tau]\cdot a_{i,j}^{t}$
11:     $\mathcal{L}_{\ominus}^{t}\leftarrow\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}\leq\tau]\cdot a_{i,j}^{t}$
12:     $\mathcal{L}_{\text{enr}}^{t}\leftarrow\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}$
13:     $\mathcal{L}_{\text{ent}}^{t}\leftarrow\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}>\tau]\cdot\mathcal{H}(\bar{A}^{t}_{j,i})$
14:     $\mathcal{L}_{\text{s}}^{t}\leftarrow\sum_{j=1}^{T^{\prime}}\|\bar{A}^{T}_{j,i}-\bar{A}^{t}_{j,i}\|_{2}^{2}$
15:     $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}_{\text{c}}^{t}+\lambda_{1}\mathcal{L}_{\text{enr}}^{t}+\lambda_{2}\mathcal{L}_{\text{ent}}^{t}+\lambda_{3}\mathcal{L}_{\text{s}}^{t}$
16:   end for
17:   $z_{t}^{\prime}\leftarrow z_{t}-\alpha_{t}\nabla_{z_{t}}(\mathcal{L}/N)$
18:   if All $\mathcal{L}_{\text{corr}}^{t}\leq\tau_{\text{corr}}$ then
19:     break
20:   end if
21:   $z_{t}\leftarrow z_{t}^{\prime}$
22:  end for
23:  $z_{t-1}\leftarrow\operatorname{DiT}(z_{t}^{\prime},\mathcal{P},t)$
24:end for
25:Continue denoising using the standard denoising diffusion process for $t=T{-}k,\dots,1$
26:return $z_{0}$
.7 Experimental Setup
Hyperparameters were tuned early using a small validation set specific to each task. While large-scale ablations were limited by computational constraints, we found the method to be robust across scenarios with only minor adjustments.
All experiments are conducted using Wan 2.1 as the backbone. We apply a classifier-free guidance scale of 6 and a sample shift of 3 during generation.
For single-object scenes, we use $\lambda_{1}{=}1$, $\lambda_{2}{=}10$, $\lambda_{3}{=}1$, a learning rate of $0.0005$, a Pearson stopping threshold of $\tau_{\text{corr}}{=}0.9$, and apply optimization for the first l=10l{=}10 denoising steps with up to k=10k{=}10 gradient updates per step.
In two-object scenes, which require more aggressive guidance, we raise the learning rate to $0.001$ for improved convergence. For motion-centric prompts, we reduce the temporal alignment weight to $\lambda_{1}{=}0.3$ and relax the Pearson threshold to $\tau_{\text{corr}}{=}0.85$ to allow more flexible optimization.
.8 Applications Setup
Single-Object Temporal Control.
To assess temporal grounding for individual objects, we construct a dataset using the 80 object categories detectable by the YOLO detector. These span common classes, such as animals and vehicles, as well as specific items, including handbags and sports equipment. Videos are generated from prompts of the form: “An empty scene. Suddenly, $\langle o_{1}\rangle$ appears out of nowhere, drawing all attention.” Object classes are uniformly distributed across different time intervals.
Two-Object Temporal Control.
We evaluate TempoControl on 82 object pairs from the VBench protocol. Prompts take the form: “The video begins with a serene view centered on $\langle o_{1}\rangle$, with no sign of $\langle o_{2}\rangle$. In the second half, $\langle o_{2}\rangle$ unexpectedly appears, altering the dynamic of the scene.” A static control signal (all ones) is used for $\langle o_{1}\rangle$. We sample 20 evenly spaced frames and require $\langle o_{1}\rangle$ to always be present. A frame is counted as successful if both objects are detected when $\langle o_{2}\rangle$ is active, and only $\langle o_{1}\rangle$ is detected when $\langle o_{2}\rangle$ is inactive.
Motion Temporal Control.
To test fine-grained temporal modulation of actions, we generate 100 videos using prompts such as: “A video of $\langle s\rangle$ $\langle v\rangle$ $\langle a\rangle$ with a strong movement at the $\langle t\rangle$ second,” where $\langle s\rangle$, $\langle v\rangle$, $\langle a\rangle$, and $\langle t\rangle$ denote subject, verb, adverb, and time reference, respectively. Example: “A child jumping joyfully at the third second.” The verb is controlled via a binary temporal mask (1 only at $\langle t\rangle$), while the subject is always present.
Multi-Object Accuracy.
We benchmark TempoControl on the VBench Multiple Objects benchmark, which evaluates the generation of all the mentioned objects. Each video is sampled at 16 evenly spaced frames. A frame is successful if both target objects are detected. The final score is the average per-frame success rate across all videos. We use all-one control signals for both objects and optimize only the first 4 denoising steps, skipping the Pearson-based stopping criterion.
.9 Human Evaluation Protocol
To assess temporal alignment and visual quality, we conducted a human preference study using side-by-side video comparisons. For each prompt, two videos were presented, one generated by our method and one by the text baseline, randomly assigned as Video A or B to avoid bias.
Annotators were asked to read the prompt and answer two questions: (1) Which video is more temporally accurate (i.e., does the object appear at the correct time and remain visible)? and (2) Which video looks more visually appealing? A “Same (Use sparingly!)” option was available if no difference was perceived.
Example screenshots of the evaluation interface are shown in Figures 6 and 7.
Figure 6: Evaluation interface showing the prompt and two videos (A and B) for comparison.
Figure 7: Second example of the human evaluation interface.
.10 Multi-Object Evaluation
We report extended results on multi-object text-to-video generation using the VC-Bench benchmark. In this setting, prompts reference multiple entities with temporal conditions (e.g., “A dog in the beginning, then a cat”). We compare our method against a text baseline.
Table 4 summarizes results. Our method achieves substantial gains in object-centric metrics. Specifically, we improve multi-object accuracy from 74.1% to 76.4% (GriT) and from 61.5% to 65.7% (YOLO), demonstrating improved temporal grounding and visibility of both entities. We also observe improvements across all consistency metrics, including subject consistency (from 97.2% to 97.8%) and background consistency (from 97.6% to 98.1%).
The dynamic degree shows a slight reduction (from 30.5% to 18.8%), which reflects a known tradeoff: our method promotes temporal consistency, often leading to more stable object appearance throughout the video. As a result, motion may be reduced.
However, we note a limitation of this metric: VC-Bench implicitly favors videos where objects are present in all frames. A more suitable benchmark might instead reward appearances rather than requiring full-video presence.
Table 4: Full results for Multiple object benchmark results. Best results per column are in bold.
Method Multiple Object (GriT) Multiple Object (YOLO) Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality Text 74.13% 61.54% 97.22% 97.56% 99.18% 30.49% 62.84% 70.25% Ours 76.37% 65.73% 97.81% 98.10% 99.40% 18.78% 62.52% 70.21%
.11 Ablations
Table 5 presents an ablation study in the two-object setting, evaluating the impact of removing the energy term, entropy term, and spatial penalty.
Omitting the energy term results in a modest decrease in temporal accuracy (from 55% to 52.56%). In contrast, removing the entropy term has a more significant effect: temporal accuracy falls to 48.96%, and the temporal presence accuracy drops sharply from 43.54% to 34.76%. Subject Consistency, Aesthetic Quality, and Imaging Quality also decline. Without entropy regularization, attention becomes overly diffuse, often blending the two objects.
Qualitative analysis shows that without the entropy term, the model frequently replaces the static object with the dynamic one (see Section .8 for details on the static and temporally conditioned objects, and Figure 8 for qualitative results).
Table 5: Ablation results for the two-object setting.
Method Temporal Accuracy T. Absence Accuracy T. Presence Accuracy Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality Text baseline 37.50% 45.85% 29.15% 92.06% 95.32% 98.92% 62.20% 53.26% 68.56% Ours 55.00% 66.46% 43.54% 92.21% 95.43% 99.19% 41.46% 56.55% 70.66% No energy 52.56% 58.90% 46.22% 92.31% 95.45% 99.34% 46.34% 56.12% 71.15% No entropy 48.96% 63.17% 34.76% 89.99% 94.39% 98.92% 53.66% 54.66% 68.95% No spatial 52.01% 60.85% 43.17% 92.76% 95.67% 99.23% 48.78% 56.75% 70.25%
Figure 8: Ablation study. Left: without entropy regularization, the motorcycle disappears. Right: the bear is incorrectly replaced by a zebra.
.12 Impact of Temporal Phrasing
We analyze the effect of using explicit temporal phrasing (e.g., “in the third second,” “in the second part of the video”) in the prompt for the one-object setup. Table 6 compares Wan 2.1 and our method, each evaluated with and without such phrasing.
For Wan 2.1, omitting temporal cues slightly improves temporal accuracy (63.94% to 65.18%) and yields higher aesthetic (45.98% to 48.78%) and imaging quality (53.76% to 59.99%), suggesting that the model struggles to interpret or apply explicit timing effectively.
Our method benefits even more: removing temporal phrases results in the highest temporal accuracy (81.12%) along with gains in aesthetic (48.16%) and imaging quality (56.32%). We hypothesize that Wan 2.1 is not trained to handle precise temporal expressions, and current attention mechanisms may not effectively capture their semantics. As a result, such phrasing introduces ambiguity. In contrast, concise prompts without explicit timing allow our method to align more reliably with the intended temporal structure.
For fairness, all main paper results use prompts with temporal phrasing to match the baseline. However, these findings suggest that our method achieves even stronger performance when using simpler prompts without explicit timing constraints.
Table 6: Effect of temporal phrasing in prompts for the one-object setup. Removing explicit timing improves video quality.
Method Temp. Acc. T. Abs. Acc. T. Pres. Acc. Subj. Cons. Backg. Cons. Motion Smooth. Dyn. Deg. Aesthetic Imaging Wan with time 63.94% 67.38% 60.50% 84.66% 92.30% 99.16% 47.50% 45.98% 53.76% Wan without time 65.18% 65.12% 65.25% 85.29% 92.44% 99.22% 45.00% 48.78% 59.99% Ours with time 81.00% 88.75% 72.91% 80.60% 90.33% 99.05% 38.75% 44.37% 52.99% Ours without time 81.12% 84.75% 77.50% 81.29% 90.72% 99.06% 35.00% 48.16% 56.32%
.13 Additional Qualitative Examples
We provide qualitative results across the one-object, two-object, and motion-based setups. For each example, we show sampled frames from the generated videos, illustrating the temporal grounding of object appearances. Prompts are abbreviated for brevity, indicating the key object and its expected time of appearance. Note that once introduced, the object is expected to persist until the end of the video. Videos can be viewed on the project page.
TextOurs
apple 4sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
bear 3sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
bench 2sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
bus 2sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
clock 5sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
elephant 3sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
surfboard 3sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
tie 3sec
TextOurs
umbrella 3sec
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
vase 5sec
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
zebra 3sec
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
wine glass 4sec
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
kite 3sec
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
fire hydrant 2sec
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
skateboard 3sec
Figure 9: One object Text vs Ours.
TextOurs
a bird and a cat
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a boat and an airplane
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a book and a clock
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a carrot and an umbrella
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a chair and a couch
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a giraffe and a bird
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a horse and a sheep
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a teddy bear and a frisbee
TextOurs
a train and a boat
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
a truck and a bicycle
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
a wine glass and a chair
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
a zebra and a giraffe
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
an umbrella and a handbag
TextOurs
a bear and a zebra
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
a cow and an elephant
Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
a cat and a dog
Figure 10: Two objects Text vs Ours.
TextOurs
chimpanze clapping 1sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
horse shaking its mane 4sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
lion roaring 1sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
magician pulling rabbit 5sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
man lifting weights 3sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
rabbit thumping 4second
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
snake slithering 1sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
tiger leaping 1sec
TextOurs
child jumping 1sec
Refer to caption Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
wolf howling 2sec
Figure 11: Movement Text vs Ours.

### v2

@@
 TempoControl utilizes cross-attention maps, a key component of text-to-video diffusion models, to guide the timing of concepts through a novel optimization approach.
-Our method steers attention using three complementary principles: aligning its temporal shape with a control signal (via correlation), amplifying it where visibility is required (via energy), and maintaining spatial focus (via entropy).
-TempoControl allows precise control over timing while ensuring high video quality and diversity.
-We demonstrate its effectiveness across various video generation applications, including temporal reordering for single and multiple objects, as well as action and audio-aligned generation.
+Our method steers attention using three complementary principles: aligning its temporal pattern with a control signal (correlation), adjusting its strength where visibility is required (magnitude), and preserving semantic consistency (entropy).
+TempoControl provides precise temporal control while maintaining high video quality and diversity.
+We demonstrate its effectiveness across various applications, including temporal reordering of single and multiple objects, action timing, and audio-aligned video generation.
+Please see our project page for more details: https://shira-schiber.github.io/TempoControl/.
 Figure 1: Applications of our inference-time temporal control method (TempoControl), which enables multi-object and single-object control, motion control, and audio alignment.
@@
 Words that appear according to a temporally varying control signal are shown in orange.
-Codehttps://github.com/Shira-Schiber/TempoControl
-Project Pagehttps://shira-schiber.github.io/TempoControl/
 1 Introduction
-Generative video models have improved dramatically, enabling the synthesis of high-quality, temporally coherent videos from natural language descriptions (Singer et al.
-2022; Blattmann et al.
-2023; Yang et al.
-2024; Bar-Tal et al.
-2024; Wan et al.
-2025).
+Generative video models have improved dramatically, enabling the synthesis of high-quality, temporally coherent videos from natural language descriptions [22, 6, 34, 3, 23].
 These models have enabled creative applications in animation, design, and virtual content creation by generating visually compelling sequences that align with textual descriptions.
@@
 This includes making an object appear in the middle of a scene or aligning visual aspects of lighting with the sound of thunder (see Figure 1).
-Recent works on generative video modeling have proposed various spatial and motion control mechanisms, including object masking, layout conditioning, or added camera controls (Gupta et al.
-2024; Wang et al.
-2024b; Feng et al.
-2025; He et al.
-2024).
+Recent works on generative video modeling have proposed various spatial and motion control mechanisms, including object masking, layout conditioning, or added camera controls [10, 25, 9, 11].
 In contrast, temporal control has received significantly less attention.
 Incorporating such control typically requires additional supervision in the form of temporally annotated datasets.
-Since video data is often scarce, extending annotations to capture temporal variation is expensive and often impractical at scale.
+Since video data is often scarce, extending annotations to capture temporal variation is expensive and usually impractical at scale.
 Synthetic data could be used; however, generating diverse and temporally accurate video-text pairs, especially for abstract concepts such as movement, is a highly non-trivial challenge.
 In this work, we propose a lightweight yet effective inference-time control method, TempoControl, which introduces temporal control for generated video.
-Instead of retraining the generative model, TempoControl leverages a property already embedded in these systems: the text-frame cross-attention mechanism.
+Instead of retraining the generative model, TempoControl leverages a property already embedded in these systems: the text-vision cross-attention mechanism.
 Cross-attention layers inherently encode strong signals about which specific words in the prompt are realized in each generated frame.
@@
 Attention control is achieved by carefully steering the latent variables during the diffusion process.
-At each denoising step, we apply a few stochastic gradient descent iterations (Battash, Wolf, and Lindenbaum 2024) until a satisfactory level of temporal alignment is achieved, notably, without updating model parameters.
-To guide temporal alignment, TempoControl introduces a novel loss with three complementary components, based on: Pearson correlation, energy, and entropy.
+At each denoising step, we apply a few stochastic gradient descent iterations [7, 5] until a satisfactory level of temporal alignment is achieved, notably, without updating model parameters.
+To guide temporal alignment, TempoControl introduces a novel loss with three complementary components, based on Pearson correlation, magnitude, and entropy.
 The Pearson term encourages the temporal shape of attention to match that of an external control signal, operating on the normalized attention vector to align the presence of a concept with the desired timing.
 However, since correlation is scale-invariant, it may yield high scores even when attention values are too low to render the object visible.
-To mitigate this, we incorporate an energy term that directly promotes stronger attention in frames where the temporal signal is high, and suppresses it elsewhere.
+To mitigate this, we incorporate a magnitude term that directly promotes stronger attention in frames where the temporal signal is high, and suppresses it elsewhere.
 Additionally, we introduce an entropy regularization term to maintain spatial focus, ensuring that when attention is activated, it remains coherent and not diffusely spread across the frame.
 Beyond single-object scenarios, as shown in Figure 1, our method generalizes well to multi-object scenes.
-In such cases, it is important to maintain the visual grounding of non-target entities; we find that preserving one object (e.g., the bird) while selectively manipulating another (e.g., the cat) helps retain the overall scene structure.
+In such cases, it is crucial to maintain the visual grounding of non-target entities.
+We find that preserving one object (e.g., the bird) while selectively manipulating another (e.g., the cat) helps retain the overall scene structure.
 In addition to nouns, our method can also manipulate verbs (e.g., clapping), effectively modifying the motion depicted in the video.
@@
 While we focus on simple cases involving a single audio event (e.g., a lightning strike), we observe promising results, suggesting that our approach may naturally extend to zero-shot multimodal control without additional training.
-Our contributions are as follows: (i) We propose an inference-time method for temporally controlling the appearance of specific words in text-to-video diffusion models, without requiring model finetuning or additional supervision.
-(ii) We introduce a suite of losses that operate over spatiotemporal attention to enable fine-grained control over both the timing and intensity of word grounding.
-(iii) We demonstrate the versatility of our method across various settings, including multi-object control, motion manipulation, and audio-to-video alignment, and introduce a new metric to evaluate the task of temporal alignment control, which is not covered by existing benchmark suites.
-Figure 2: Video generated for the prompt “A cat and a dog.” Before (left) and after optimization (right).
-Top: Spatial attention maps for the tokens cat and dog at denoising step t=3t{=}3, shown for frames j=2,10,19j{=}2,10,19.
-Attention is overlaid on generated frames.
-Bottom: temporal attention $a_{i,j}^{t}$ (blue) vs.
-target mask $m_{i,j}$ (orange), with Pearson correlation loss shown.
+Our contributions are threefold: (i) We introduce the first inference-time method for temporally controlling the appearance of specific words in text-to-video diffusion models, without additional supervision or model finetuning, allowing temporal control while preserving the fidelity and diversity of existing models.
+(ii) We propose a set of spatiotemporal attention losses that shape both the timing and strength of word grounding, together with an entropy-based regularization term that maintains semantic consistency.
+(iii) We demonstrate broad applicability across multi-object control, motion timing, and audio-to-video alignment, and introduce a new metric for temporal alignment, a dimension not covered by current benchmarks.
+Figure 2: Motivation for our approach.
+Top: We show a video generated for the prompt “The video begins with a serene view centered on the cat, with no sign of the dog.
+In the second half, the dog unexpectedly appears, altering the dynamic of the scene.” The top row displays the attention maps for the tokens cat and dog, extracted from the denoising step t=3t{=}3, for frames j=2,6,10,14,19j{=}2,6,10,14,19.
+On the left, the video is generated without our optimization.
+Despite the prompt specifying that the dog should appear in the second half, it appears early.
+This behavior is common, as Wan 2.1 often fails to depict objects or movements according to temporal cues in the prompt.
+On the right, after applying our conditioning method, the dog correctly appears in the second half of the video.
+Bottom: Temporal attention $a_{i,j}^{t}$ (blue) versus target mask $m_{i,j}$ (orange), with the corresponding Pearson correlation loss shown.
 2 Related Work
-Motion Control.
-Fine-grained motion control has been an active area of research.
-Boximator (Wang et al.
-2024b) introduces control by allowing users to specify the spatial locations of a start box and an end box for each object.
-This control is implemented by fine-tuning a dedicated control module.
-BlobGEN-Vid (Feng et al.
-2025) enhances control by introducing visual primitive-based guidance, such as lines, to enable detailed manipulation of object motion, fine-grained appearance, and smooth object transitions.
-CameraCtrl (He et al.
-2024) addresses controllability by enabling precise camera movement and perspective control.
-This is accomplished by training a plug-and-play camera pose control module on top of a video diffusion model.
-Motion Prompting (Geng et al.
-2025) enables movement control by incorporating spatio-temporally sparse or dense motion trajectory inputs.
-However, these approaches require extensive datasets and computational resources for training, substantial user annotations during inference, and notably lack precise control over motion timing or peak movement appearance.
 Attention-based Control.
-Attention-based methods have improved spatial control in text-to-image generation.
-Prompt-to-Prompt (Hertz et al.
-2023) enables word-level edits via cross-attention manipulation, while Attend-and-Excite (Chefer et al.
-2023) ensures all mentioned objects are depicted by balancing token attention.
-Battash et al.
-(2024) introduced a layout optimization strategy for multi-subject generation in text-to-image diffusion, which constrains cross-attention to improve the spatial separation of objects.
-Be Yourself (Dahary et al.
-2024) uses bounded attention to preserve subject identity and prevent semantic leakage in multi-subject text-to-image generation.
-Stable Flow (Avrahami et al.
-2025) enables training-free image editing by selectively injecting attention features into vital layers of a diffusion transformer.
-These methods operate on static images, whereas we extend attention-based control to the temporal domain, enabling control over object timing and motion in video generation.
+Attention manipulation has become a powerful tool for controlling semantics in text-to-image generation.
+Methods such as Attend-and-Excite [7, 17, 4, 8] perform inference-time optimization over cross-attention maps to ensure that all textual entities are faithfully represented in the generated image.
+Stable Flow [1] further introduces a training-free editing framework that selectively injects attention features into key layers of a diffusion transformer.
+Additional approaches, including Prompt-to-Prompt [12], leverage attention restructuring for fine-grained, text-driven edits.
 Inference time optimization.
 Inference-time optimization for text-to-video diffusion has only recently attracted attention.
-(Li et al.
-2025) pre-compute a spatio-temporal sketch and inject plan-consistent noise during sampling to tighten prompt alignment.
-(Wei et al.
-2025) converts a single keyframe into a textured, deformable 3D mesh that is re-rendered at every denoising step to maintain consistency in the garment’s appearance.
-(Shaulov et al.
-2025) derive variance-based optical-flow cues from the model’s latent predictions and add a flow-coherence loss on-the-fly, smoothing motion without retraining.
-(Lian et al.
-2023) let a large language model generate scene graphs and temporal layouts that act as attention masks, improving complex spatio-temporal grounding.
+[16] pre-compute a spatio-temporal sketch and inject plan-consistent noise during sampling to tighten prompt alignment.
+[28] converts a single keyframe into a textured, deformable 3D mesh that is re-rendered at every denoising step to maintain consistency in the garment’s appearance.
+[21] derive variance-based optical-flow cues from the model’s latent predictions and add a flow-coherence loss on-the-fly, smoothing motion without retraining.
+[18] let a large language model generate scene graphs and temporal layouts that act as attention masks, improving complex spatio-temporal grounding.
 While these methods improve fidelity and consistency without necessitating additional training, none offer explicit, user-defined control over the timing of textual element unfolding.
+Controllable Video Generation Controllable video generation has advanced rapidly, with methods exploring a broad set of control modalities.
+Early multimodal systems, such as VideoComposer [26], incorporate depth, sketch, and motion vector guidance to steer video synthesis.
+More recent approaches introduce explicit camera and object motion control on top of text-to-video generation [27, 11, 15, 33, 9, 2].
+Another line of work investigates the controllability of image-to-video object motion [25, 32].
+While effective, these training approaches rely on large paired datasets and extensive supervision, which can be costly and limit generalization.
+A parallel line of work studies training-free controllability, including CamTrol [13] for camera motion, MotionClone and DiTFlow [19, 20] for transferring motion from reference videos, and MOFT [31] for editing motion at inference.
+These rely on dense motion trajectories, but cannot specify when a concept should appear or disappear.
+The most related work, MinT [30], arranges multiple textual instructions in temporal order by learning when each should begin.
+However, MinT provides only coarse event sequencing, whereas our method offers fine-grained timing for individual concepts.
+Moreover, MinT requires fine-tuning on temporally annotated caption datasets centered on human activities, which limits its generality across diverse prompts.
+Our approach operates entirely at inference time and generalizes across settings without any training or domain-specific supervision.
 3 Preliminaries
-We build TempoControl on top of Wan 2.1 (Wan et al.
-2025), a state-of-the-art diffusion model that serves as our backbone for text-to-video generation.
+We build TempoControl on top of Wan 2.1 [23], a state-of-the-art diffusion model that serves as our backbone for text-to-video generation.
 Let $v\in\mathbb{R}^{T_{v}\times H_{v}\times W_{v}\times 3}$ represent the input video, where $T_{v}$ is the number of frames, $H_{v}\times W_{v}$ denotes the spatial resolution of each frame, and the final dimension corresponds to the three RGB color channels.
-The video is first encoded by a 3D causal variational autoencoder into a latent tensor
+A 3D causal variational autoencoder first encodes the video into a latent tensor
 $z\in\mathbb{R}^{T^{\prime}\times H^{\prime}\times W^{\prime}\times C},$ (1)
@@
 At each timestep $t$, a transformer-based network $\operatorname{DiT}(z_{t},\mathcal{P},t)$ predicts the noise in $z_{t}$, conditioned on the prompt $\mathcal{P}$ and the timestep $t$.
-3.1 Cross-attention in text-to-video diffusion models
-A key challenge in text-to-video generation is ensuring that the synthesized video aligns semantically and temporally with the input prompt.
+3.1 Cross-attention in video diffusion models
+A key challenge in text-to-video generation is maintaining semantic and temporal alignment with the input prompt.
 In diffusion-based architectures, this alignment is achieved through cross-attention layers within the transformer blocks at each denoising step $t$.
@@
 The vector $\bar{A}^{t}_{i}$ can be reshaped at temporal index $j$, yielding $\bar{A}^{t}_{j,i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$, which represents the spatial influence of word $i$ at temporal position $j$ and denoising step $t$.
-Figure 2 illustrates this for the words cat and dog in the prompt “A cat and a dog.” The attention maps are up-sampled to the frame resolution for visualization.
+Figure 2 illustrates this for the words cat and dog in the prompt “A cat and a dog.
+In the second half of the video, the dog appears.” Attention maps are upsampled to frame resolution for display.
 The attention map captures the spatiotemporal grounding of words in the generated video; TempoControl is specifically designed to steer the temporal influence of a token.
-Therefore, we focus on the scalar attention value $\hat{A}^{t}_{j,i}=\langle\bar{A}^{t}_{j,i}\rangle_{x,y}\in\mathbb{R}$, where $\langle\cdot\rangle_{x,y}$ denotes summation over the spatial coordinates $x\in[1,H^{\prime}]$ and $y\in[1,W^{\prime}]$.
-Intuitively, $\hat{A}^{t}_{j,i}$ measures the influence of word $i$ on the latent representation at temporal position $j$ during the denoising step $t$.
-To summarize the temporal attention pattern of word $i$ at denoising step $t$, we collect the scalar values across frames into a vector:
+Thus, we focus on the scalar attention value $\hat{A}^{t}_{j,i}=\langle\bar{A}^{t}_{j,i}\rangle_{x,y}\in\mathbb{R}$, where $\langle\cdot\rangle_{x,y}$ denotes summation over the spatial coordinates $x\in[1,H^{\prime}]$ and $y\in[1,W^{\prime}]$.
+Intuitively, $\hat{A}^{t}_{j,i}$ measures the influence of word $i$ on the latent representation at temporal position $j$ during denoising step $t$.
+To aggregate the temporal attention pattern of word $i$ at denoising step $t$, we collect the scalar values across frames into a vector:
 $a_{i}^{t}=[\hat{A}^{t}_{1,i},\hat{A}^{t}_{2,i},\ldots,\hat{A}^{t}_{T^{\prime},i}]\in\mathbb{R}^{T^{\prime}}.$ (2)
-We omit the first frame due to unstable attention values that do not reliably reflect word grounding. This vector serves as the target for our steering method, described in the next section.
+We omit the first frame due to unstable attention values that do not reliably reflect word grounding. This vector serves as the target for our steering method, described next.
 Figure 3: Illustration of TempoControl.
@@
 4 Method
-Given a video $v$, our goal is to generate an edited version $v^{\prime}$ such that specific tokens appear according to a temporal condition defined by a weighting mask $m_{i}=[m_{i,1},\ldots,m_{i,T^{\prime}}]$, where $m_{i,j}\in[0,1]$ indicates the desired strength of activation of token $p_{i}$ at temporal position $j$ (see Figure 2 bottom).
-In most experiments, we use binary masks (i.e., values of zero or one), except for the audio-alignment setting, where the mask reflects sound strength.
-The main steps of our approach are detailed in the appendix Algorithm 1, illustrated in Figure 3, and discussed below.
-We steer the latent representation during the first $k$ denoising steps by performing $l$ gradient update iterations at each step using the AdamW optimizer.
-For clarity, the update can be expressed as:
+Given a video $v$, our objective is to generate an edited version $v^{\prime}$ such that specific tokens appear according to a temporal condition.
+To achieve fine-grained temporal control, we employ masks, which provide a more precise and less ambiguous mechanism than textual descriptions.
+We denote as input a binary mask $m_{i}=[m_{i,1},\ldots,m_{i,T^{\prime}}]$, where each element $m_{i,j}\in\{0,1\}$ indicates whether token $p_{i}$ should appear at temporal position $j$ (see Figure 2, bottom, orange line).
+In this example, the mask specifies that the token dog should appear only in the second half of the video (i.e., frames 10–20).
+In general, $m_{i,j}$ may also take continuous values to represent the strength of appearance.
+In most experiments, we use binary masks, except in the audio-alignment setting, where the mask encodes sound intensity.
+The main steps of our approach are illustrated in Figure 3 and discussed below.
+We also provide an algorithmic description in the appendix.
+We steer the latent representation during the first $k$ denoising steps by applying $l$ gradient-based updates at each step, optimized with AdamW.
+Let $z_{t}$ denote the latent at step $t$; the update is:
 $z^{\prime}_{t}=z_{t}-\alpha\nabla_{z_{t}}\mathcal{L}^{t},$ (3)
-where $\alpha\in\mathbb{R}$ is the learning rate and $t$ is the current denoising step.
-The loss $\mathcal{L}^{t}$ consists of four components (detailed below): (i) temporal correlation, (ii) energy, (iii) entropy regularization, and (iv) spatial consistency.
+where $\alpha$ is the learning rate.
+The loss $\mathcal{L}^{t}$ consists of a single main component: a temporal correlation term that aligns the attention pattern with the target mask.
+While this term is most significant for aligning with the temporal condition, we also include two terms that further improve the results: (ii) a magnitude term that amplifies or suppresses token-level activation, and (iii) an attention entropy regularization term that preserves spatial consistency.
+We describe each component in detail below.
 Temporal correlation term.
-The first loss term encourages the temporal attention vector $a_{i}^{t}$ to match the target mask $m_{i}$ using the Pearson correlation:
+The main loss term encourages the temporal attention vector $a_{i}^{t}$ to match the target mask $m_{i}$ using the Pearson correlation:
 $\mathcal{L}_{\text{corr}}^{t}=-\frac{\text{Cov}(m_{i},\tilde{a}_{i}^{t})}{\sigma_{m_{i}}\sigma_{\tilde{a}_{i}^{t}}},$ (4)
-where $\tilde{a}_{i}^{t}$ is the min-max normalized version of $a_{i}^{t}$, defined as $\tilde{a}_{i}^{t}=\frac{a_{i}^{t}-\min(a_{i}^{t})}{\max(a_{i}^{t})-\min(a_{i}^{t})}$.
-Here, $\text{Cov}(\cdot,\cdot)$ denotes empirical covariance, and $\sigma_{m_{i}}$ and $\sigma_{\tilde{a}_{i}^{t}}$ are the corresponding standard deviations.
-Note that if $m_{i}$ has zero variance, such as when encouraging an object to appear throughout the entire video, the correlation term is undefined; in such cases, it is excluded from the loss.
-This alignment term adjusts the temporal placement of words in the video but does not affect their overall strength.
-As a result, words that are initially invisible may remain so.
-The following loss term is designed to modulate the temporal strength of attention.
-Attention energy term.
-The energy term adjusts the overall magnitude of attention for word token $i$ based on the temporal mask $m_{i}$.
-We use the indicator function $\mathbb{1}_{\{\cdot\}}$, which equals 1 when the condition is true and 0 otherwise.
-The positive energy encourages attention where the word should be active:
+where $\tilde{a}_{i}^{t}$ is the min–max normalized version of $a_{i}^{t}$, defined as $\tilde{a}_{i}^{t}=\frac{a_{i}^{t}-\min(a_{i}^{t})}{\max(a_{i}^{t})-\min(a_{i}^{t})}$.
+Here, $\text{Cov}(\cdot,\cdot)$ denotes the empirical covariance, and $\sigma_{m_{i}}$ and $\sigma_{\tilde{a}_{i}^{t}}$ are the corresponding standard deviations.
+Note that if $m_{i}$ has zero variance, such as when encouraging an object to appear throughout the entire video, the correlation term becomes undefined; in such cases, it is omitted from the loss.
+The Pearson correlation measures the linear relationship between the mask $m_{i}$ and the attention term $a_{i}^{t}$, but it does not capture their scale.
+In our experiments, we observe two phenomena: (i) words that are initially invisible may remain so, and (ii) completely suppressing the appearance of a token is difficult.
+The following loss term is therefore designed to modulate the temporal strength of attention.
+Attention magnitude term.
+The magnitude term adjusts the overall strength of attention for each word token $i$ according to the temporal mask $m_{i}$, complementing the correlation-based term by restoring scale sensitivity.
+We employ the indicator function $\mathbb{1}_{\{\cdot\}}$, which equals one when the specified condition holds and zero otherwise.
+The positive magnitude component encourages attention in regions where the word is expected to be active:
 $\mathcal{L}_{\oplus}^{t}=\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}_{\{m_{i,j}>\tau\}}\cdot a_{i,j}^{t}.$ (5)
-The negative energy penalizes attention where the word should not be visible:
+The negative magnitude component penalizes attention where the word should not be visible:
 $\mathcal{L}_{\ominus}^{t}=\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}_{\{m_{i,j}\leq\tau\}}\cdot a_{i,j}^{t}.$ (6)
-The final energy loss is defined as the net difference:
-$\mathcal{L}_{\text{energy}}^{t}=\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}.$ (7)
+The final magnitude loss is defined as the net difference:
+$\mathcal{L}_{\text{mag}}^{t}=\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}.$ (7)
+Figure 4: Entropy regularization helps preserve semantic fidelity.
 Here, $\tau\in[0,1]$ is a threshold that determines whether a given time index is considered active for token $p_{i}$.
 While $m_{i}$ is often binary (i.e., 0 or 1), it may also take continuous values (e.g., for soft audio-driven conditioning).
-In such cases, the indicator $\mathbb{1}_{\{m_{i,j}>\tau\}}$ selects the time-steps where the target activation is sufficiently strong.
-Increasing the attention magnitude without constraint can lead to overly diffuse spatial maps (Qi et al.
-2025), especially in multi-object scenes, which degrades the visual grounding of word tokens.
-To regularize this behavior and preserve spatial consistency, we introduce additional loss terms in the following paragraph.
-Entropy regularization.
-Word-level attention is typically spatially focused rather than dispersed (see Figure 2).
-To discourage excessive spreading of attention during optimization, we regularize the entropy of the spatial attention maps.
-Specifically, we apply this regularization only at temporal indices where the word is expected to be active (i.e., $m_{i,j}>\tau$), resulting in:
+In such cases, the indicator $\mathbb{1}_{\{m_{i,j}>\tau\}}$ selects time steps where the target activation is sufficiently strong.
+Both the correlation and magnitude terms target the temporal objective but can harm spatial consistency.
+In our experiments, these terms sometimes corrupt object semantics, such as a smartphone turning into a distorted two-screen tablet (Figure 4; more examples in the appendix).
+To address this and preserve spatial coherence, we introduce an entropy-based regularization term.
+Attention entropy regularization.
+Optimizing for the temporal objective may compromise spatial properties, often resulting in overly spread attention.
+We counter this by regularizing the entropy of the spatial attention maps at temporal indices where the word is active (i.e., $m_{i,j}>\tau$), resulting in:
 $\mathcal{L}_{\text{entropy}}^{t}=\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}>\tau]\cdot\mathcal{H}(\bar{A}^{t}_{j,i}).$ (8)
 Here, $\mathcal{H}(\bar{A})$ denotes the Shannon entropy computed over the spatial dimensions.
-Spatial consistency penalty.
-To maintain spatial consistency, we penalize deviations between the initial and optimized spatial attention maps for the target word.
-Here, $\hat{A}^{t}_{j,i}$ denotes the spatial attention map before optimization at step $t$, for word $i$ at frame $j$.
-The penalty is defined as
-$\mathcal{L}_{\text{spatial}}^{t}=\left\|\sum_{j=1}^{T^{\prime}}\left(\hat{A}^{t}_{j,i}-\bar{A}^{t}_{j,i}\right)\right\|_{2}^{2},$ (9)
-where the difference is computed over the spatial dimensions, after summing across all latent frames.
+In our experiments, we find that discouraging excessive spreading of attention not only improves temporal presence but also enhances overall image quality (see ablations), suggesting that this term may be beneficial for video generation more broadly.
 Full objective.
 The combined loss at step $t$ is:
-$\mathcal{L}^{t}=\mathcal{L}_{\text{corr}}^{t}+\lambda_{1}\mathcal{L}_{\text{energy}}^{t}+\lambda_{2}\mathcal{L}_{\text{entropy}}^{t}+\lambda_{3}\mathcal{L}_{\text{spatial}}^{t},$ (10)
-where $\lambda_{1},\lambda_{2},\lambda_{3}$ are hyperparameters used to calibrate the contribution of the regularization terms.
-This objective is optimized over latent codes during inference; model weights remain fixed.
+$\mathcal{L}^{t}=\mathcal{L}_{\text{corr}}^{t}+\lambda_{1}\mathcal{L}_{\text{magnitude}}^{t}+\lambda_{2}\mathcal{L}_{\text{entropy}}^{t},$ (9)
+where $\lambda_{1}$ and $\lambda_{2}$ control the auxiliary terms.
+The correlation term is sufficient for most cases and provides strong temporal alignment.
+The magnitude term is helpful when the base model fails to accurately depict a token or when the token must be removed entirely.
+The entropy term maintains spatial focus and enhances video quality.
+The objective is optimized over latents during inference, while the model weights remain fixed.
 Correlation-based early stopping.
@@
 If the correlation exceeds a threshold $\tau_{\text{corr}}$, we terminate optimization early for that step; otherwise, we continue refining until the limit is reached.
-Method Temporal Accuracy T.
-Absence Accuracy T.
-Presence Accuracy Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality One Object Text 63.94% 67.38% 60.50% 84.66% 92.30% 99.16% 47.50% 45.98% 53.76% Ours 81.00% 88.75% 72.91% 80.60% 90.33% 99.05% 38.75% 44.37% 52.99% Two Objects Text 37.50% 45.85% 29.15% 92.06% 95.32% 98.92% 62.20% 53.26% 68.56% Ours 55.00% 66.46% 43.54% 92.21% 95.43% 99.19% 41.46% 56.55% 70.66% Movement Text 19% 93.17% 95.91% 98.10% 73.00% 55.81% 60.46% Ours 53% 92.86% 95.67% 98.20% 80.00% 55.06% 61.76%
-Table 1: Comparison against Wan 2.1 across three temporal-control applications: (i) one object, (ii) two objects, and (iii) movement.
-Results for irrelevant metrics are marked with –.
-Higher values indicate better performance for all metrics.
-Method Multiple Object (GriT) Multiple Object (YOLO) Dynamic Degree Text 74.13% 61.54% 30.49% Ours 76.37% 65.73% 18.78%
-Table 2: Multiple object benchmark assessing the ability to depict multiple objects across the video.
-Question Type Ours Baseline Same Temporal Accuracy 61.51% 16.94% 21.55% Visual Quality 62.66% 25.99% 11.35%
-Table 3: User study results for temporal accuracy and visual quality.
 5 Experiments
-Below, we describe the applications and evaluation protocols used in our experiments.
-Setup.
-We use Wan 2.1 1.3B as our text-to-video backbone.
-All experiments were run on a single NVIDIA H200 GPU.
-TempoControl takes approximately $\sim$460s per video, compared to $\sim$170s for vanilla generation.
-Hyperparameter details and ablations are available in the appendix.
+Experimental setup.
+All experiments are conducted on a single NVIDIA H200 GPU using Wan 2.1-S as the backbone.
+Due to limited resources, we focus on the smaller model, which already outperforms larger baselines.
+We optimize the latent representation during the first 10 denoising steps, applying up to 10 gradient updates per step with a learning rate of $5\times 10^{-4}$, increased to $1\times 10^{-3}$ for the two-object setting.
+We use $\lambda_{1}{=}0.3$ and $\lambda_{2}{=}10$.
+See the appendix for ablation details.
 Baseline.
-We employ Wan 2.1 1.3B with explicit temporal cues in the prompt (e.g., “…at the third second…”). We describe the exact phrasing details in the appendix.
+We evaluate our method on several text-to-video backbones, including Wan 2.1 (1.3B and 14B), Wan 2.2-TI2V (5B), and CogX1.5 (5B).
+As a baseline, we use each model with explicit temporal cues in the prompt (e.g., “…at the third second …”).
+The exact phrasing templates are provided in the appendix.
+Note that we also use the same timing templates in our approach for fair comparison.
+The lack of dedicated baselines stems from the novelty of our objective, which involves fine-grained temporal control and necessitates the definition of new evaluation metrics, discussed next.
 Metrics.
-We evaluate on VBench (Huang et al.
-2024), reporting temporal-related metrics: Subject Consistency (main subject coherence across frames), Background Consistency (background stability), Motion Smoothness (absence of temporal flicker), and Dynamic Degree (overall motion magnitude), as well as quality metrics: Aesthetic Quality (visual appeal) and Imaging Quality (clarity of frames).
-Additional details about these metrics are provided in the appendix.
-We also introduce a new metric, Temporal Accuracy, which measures whether objects appear in the correct frames as specified by the prompt, and provides breakdowns for both object absence and presence.
-By default, the official VBench metric uses the GRiT detector (Wu et al.
-2024).
-To provide a more up-to-date assessment, we report temporal accuracy results using YOLOv10 (Wang et al.
-2024a).
-For movement, Temporal Accuracy is assessed by matching the movement time specified in the prompt with the second exhibiting the most movement, as determined by optical flow.
-Each video is sampled at 4 frames per second; we compute the average flow magnitude per second, and the second with the highest average is considered the most active.
+Current benchmarks such as VBench [14] do not measure temporal control.
+To address this, we introduce Temporal Accuracy, which checks whether each token appears in the correct frames specified by the prompt, with separate scores for presence and absence.
+VBench relies on the GRiT detector [29], which we find unreliable for small or partially visible objects.
+Instead, we detect objects using the more recent YOLOv10 [24].
+For motion-centric prompts, Temporal Accuracy aligns the prompt-specified action time with the second exhibiting the strongest motion, as determined by per-second optical flow magnitude computed from videos sampled at 4 fps.
+Full details are provided in the appendix.
+We also report results on the VBench Multiple Object benchmark (detecting object appearance throughout the video), and the Imaging Quality metric that uses a perceptual image-quality model.
+Benchmark.
+In the single-object setup, we utilize 80 YOLOv10 classes, which span animals, everyday objects, vehicles, and more.
+For the two-object setting, we use 82 object pairs based on the VBench multiple-object benchmark.
+For movement, we curate 100 types of actions.
+Temporal masks are generated together with the prompt templates to ensure alignment.
+Full prompt patterns and categories are provided in the appendix.
+Table 1: Comparison across one-object, two-object, and movement temporal-control tasks.
+Baselines use text-only temporal cues, while Ours applies TempoControl on Wan2.1 (1.3B).
+Bold values indicate the best result in each block.
+Method # Temporal Accuracy T. Absence Accuracy T. Presence Accuracy Imaging Quality
+One Object
+CogX1.5 5B 61.43% 48.64% 75.50% 65.36%
+Wan2.1-S 1.3B 63.94% 67.38% 60.50% 53.76%
+Wan2.1-L 14B 64.23% 55.34% 74.00% 66.89%
+Wan2.2 5B 78.02% 87.42% 68.00% 48.13%
+Ours 1.3B 82.50% 83.25% 81.75% 56.51%
+Two Objects
+CogX1.5 5B 24.68% 46.46% 4.88% 63.78%
+Wan2.1-S 1.3B 37.50% 45.85% 29.15% 68.56%
+Wan2.1-L 14B 40.07% 54.15% 27.27% 69.01%
+Wan2.2 5B 47.05% 56.86% 36.59% 67.83%
+Ours 1.3B 52.93% 58.05% 47.80% 71.09%
+Movement
+CogX1.5 5B 21% 61.39%
+Wan2.1-S 1.3B 19% 60.46%
+Wan2.1-L 14B 18% 61.31%
+Wan2.2 5B 23% 60.61%
+Ours 1.3B 57% 63.38%
 5.1 Quantitative Results
 Single object temporal control.
-In Table 1 (One Object), our method improves temporal accuracy by 17.06% to 81.00%, with similar gains in temporal absence (+21.37) and presence (+12.41) accuracy compared to text prompts with explicit time cues.
-Image Quality and Aesthetic Quality remain comparable, with a difference of 1%.
-We observe decreases in Dynamic Degree, dropping by 8.75% to 38.75%, reflecting reduced dynamics due to temporal control, which requires object presence over extended periods of the video.
+In Table 1 (One Object), our method improves temporal accuracy by 18.56%, with gains of 15.87% in absence accuracy and 21.25% in presence accuracy over the Wan2.1-S baseline.
+Imaging Quality also increases by about 3%.
+Larger baseline models achieve higher imaging quality, as expected, yet all fail to provide reliable temporal control from text alone.
+Wan2.2’s strong absence accuracy largely reflects degraded image quality, which causes detectors to miss objects.
+In a separate experiment, we tested removing the timing cues, which increased its Imaging Quality to 57.78%, indicating that it is not robust to temporal instructions; details are in the appendix.
 Two objects temporal control.
-In Table 1 (Two Objects), we show that the two-object setting presents a significant challenge, with the baseline achieving only 37.5% temporal accuracy.
-Our method improves to 55.0%.
-Temporal absence and presence accuracies also show gains of 20.6% and 14.4%, respectively.
-Imaging and Aesthetic Quality increase by about 2% each.
-Subject Consistency, Background Consistency, and Motion Smoothness remain similar between methods.
-At the same time, we find that Dynamic Degree is reduced here, as in the one-object case, by roughly 21%.
-This reflects the trade-off imposed by temporal control, which requires static background objects and reduces dynamic content.
-Figure 4: Examples of videos generated by Wan 2.1 (left) and our method with temporal conditioning (right) for prompts that specify temporal conditions.
-The prompts are simplified for brevity.
-We show frames corresponding to the first, middle, and last third of each video.
-Figure 5: Examples of video alignment to an audio signal.
+In Table 1 (Two Objects), the two-object setting is notably challenging.
+Wan2.1-S baseline achieves only 37.50% temporal accuracy, while our method improves this to 52.93% (+15.43).
+Absence and presence accuracies also increase by 12.20% and 18.65%, respectively, and Imaging Quality improves by 2.53%.
+Larger models also show similar degradation in temporal control in the two-object setting.
+Conversely, Imaging Quality is consistently higher across all methods, suggesting a possible bias in the perceptual metric toward multi-object scenes.
+Our method achieves the highest Imaging Quality.
 Movement temporal control.
-Table 1 (Movement) shows that TempoControl significantly improves temporal accuracy (53% vs.
-19%) compared to the text-prompt baseline.
-We note that the original temporal accuracy of the model is similar to random, indicating that Wan 2.1 does not capture movement cues at all.
-The Dynamic Degree here increases, as the goal is to introduce more movement, demonstrating that our approach successfully creates more pronounced motion.
-Imaging and Aesthetic Quality remain on par with the baseline.
-Overall, our explicit temporal guidance sharpens movement timing without compromising visual quality and consistency.
+Table 1 (Movement) evaluates temporal accuracy when conditioning is indicative of a peak of action.
+The temporal accuracy measures the timing of when the action peaks (see Metrics).
+Wan2.1-S baseline achieves only 19% temporal accuracy, indicating that the model does not follow action timing from text alone.
+TempoControl improves this to 57%, showing a substantial gain in aligning motion onset with the prompt.
+Low performance in movement conditioning is consistent across all models. Our Imaging Quality remains the best in this setup as well.
 Multiple objects benchmark.
@@
 We observe improved performance on the Multiple Object metric (76.34% vs.
-74.13%), although this comes at the cost of reduced video dynamics.
+74.13%).
 Detailed results are provided in the appendix.
+Ablation.
+Table 3 shows that each component of our objective contributes differently to temporal control and visual quality.
+Pearson correlation (C) provides the strongest temporal alignment, driven largely by improvements in T.
+Absence Accuracy, but it also reduces Imaging Quality.
+Closer inspection suggests that degraded visuals can cause detectors to miss objects, artificially inflating absence scores.
+Entropy regularization (E) improves presence accuracy and stabilizes spatial attention.
+Combining both (C+E) yields balanced gains across metrics.
+Notably, using entropy alone produces the highest Imaging Quality (59.52%), outperforming even the text-only baseline, suggesting that entropy-based attention smoothing is broadly beneficial for video generation.
+Additional ablations are provided in the appendix.
 Human evaluation.
-We conducted a human evaluation with 50 annotators, each comparing 16 videos (8 one-object, 8 two-object) generated by TempoControl and Wan 2.1.
-Annotators selected which video was better in terms of temporal accuracy and visual quality, or if they were the same.
-As shown in Table 3, our method outperformed the baseline in temporal accuracy (61.51% vs.
-16.94%) and also improved visual quality.
-For temporal accuracy, 21.55% of responses indicated no preference, likely due to subtle timing differences.
-Full details of the evaluation are provided in the appendix.
+We conducted a user study with 50 computer science graduate students, each comparing 16 video pairs (8 one-object and 8 two-object) generated by TempoControl and Wan 2.1 with text-based temporal cues.
+Annotators judged temporal accuracy and visual quality, or indicated no difference, which we asked them to use sparingly (full protocol details are provided in the appendix).
+As shown in Table 4, our method is preferred for temporal accuracy (61.51% vs.
+16.94%) and also improves visual quality.
+Notably, 21.55% of temporal comparisons reported no preference, reflecting that humans struggle to perceive minor timing differences.
+Table 2: Full results for the multiple-object benchmark. Best results per column are in bold.
+Method Multiple Object (GRiT) Multiple Object (YOLO) Imaging Quality
+Text 74.13% 61.54% 70.25%
+Ours 76.37% 65.73% 70.21%
+Table 3: Ablation results for the one-object setting. C = Pearson correlation, E = Entropy.
+Method Temporal Accuracy T. Absence Accuracy T. Presence Accuracy Imaging Quality
+Text 63.94% 67.38% 60.50% 53.76%
+Only C 81.19% 91.50% 70.88% 50.96%
+Only E 72.94% 66.88% 79.00% 59.52%
+C + E 78.38% 77.38% 79.38% 57.60%
+Ours 82.50% 83.25% 81.75% 56.51%
+Table 4: User study comparing our method with Wan2.1 in terms of temporal accuracy and visual quality.
+Question Type Ours Baseline Same
+Temporal Accuracy 61.51% 16.94% 21.55%
+Visual Quality 62.66% 25.99% 11.35%
+Figure 5: Examples generated by Wan 2.1 (left) and our temporally conditioned method (right) for prompts with timing constraints.
+Prompts are simplified for brevity.
+Shown are frames from the first, middle, and final thirds of each video.
+Figure 6: Examples of video alignment to an audio signal.
 5.2 Qualitative Results
-In Figure 4, we show four videos generated to depict objects or actions at specific times.
-In the first row, where a single apple is expected to appear at the fourth second, Wan 2.1 generates the apple near the middle of the five-second video, while our method correctly places it near the end of the video.
-For the “cat and dog, dog at first half” prompt, Wan 2.1 fails to show the dog at the beginning of the video, whereas our approach accurately places the dog in the first half.
-In the “magician pulling a rabbit last second” scenario, Wan 2.1 displays the rabbit from the start, instead of only in the final second as required.
-Similarly, for “tiger leaping first second,” Wan 2.1 does not capture the leaping motion at the correct time, while our method, with temporal conditioning, accurately generates the requested action.
-All videos are included in the supplementary material.
+In Figure 5, each row illustrates a different temporal-control task.
+The first row shows a one-object example: for the prompt “an apple appears at the fourth second,” Wan 2.1 places the apple near the middle of the five-second video, whereas our method correctly places it near the end.
+The second row shows a two-object example: for “a car and a bicycle, car in the second half,” Wan 2.1 displays the car at the beginning instead of the specified time.
+The third and fourth rows depict movement-based prompts.
+For a magician pulling a rabbit at the last second, Wan 2.1 shows the rabbit throughout the video, whereas our method reveals it only at the end.
+For a tiger leaping in the first second, Wan 2.1 fails to time the leap, whereas our method accurately aligns the motion with the prompt.
+Additional examples are included in the supplementary material.
 5.3 Audio-Visual Alignment
-We also experimented with zero-shot alignment of video to external audio by using a preprocessed audio envelope as the temporal condition.
-Given a waveform $y$, we extract its onset strength envelope $s_{t}$, adaptively pool it to match the number of frames, and normalize to $[0,1]$ via min-max normalization.
-Prominent peaks are preserved via thresholding and Gaussian smoothing:
-$\tilde{s}_{t}=\begin{cases}\hat{s}_{t},&\text{if }\hat{s}_{t}\geq\tau_{audio},\\ \text{GaussianSmooth}(\hat{s}_{t};\sigma),&\text{otherwise}.\end{cases}$ (11)
-where $\tau_{audio}$ is a threshold for peak preservation, and $\sigma$ is the standard deviation of the Gaussian kernel used for smoothing.
-In Figure 5, we present examples of accurate video adaptation to audio peaks, with additional results in the appendix.
-Although experiments are limited to simple cases with prominent peaks, these findings suggest a promising direction for aligning generated video with audio without requiring costly training on curated datasets.
-5.4 Discussions
-Text temporal control limitations.
-In our experiments, adding explicit timing to Wan 2.1 prompts did not improve temporal accuracy and often reduced video and image quality.
-For example, temporal accuracy fell from 65% to 64%, and video quality from 56% to 53%.
-For fairness, all reported results use the same prompt setup as the text baseline, despite omitting explicit timing, which yields higher performance for our approach.
-For instance, for a single object, we observe a video quality of 56% compared to 53%, with the same temporal accuracy of 81%.
-See the appendix for details.
-Consistency limitations.
-Temporal conditioning may lead to perceptual changes in background or unspecified features.
-For example, it can alter the color of a cat (see Figure 1).
-Although perceptual consistency is often preserved, our current objectives do not fully guarantee it, which remains a promising area for future work.
+We also explore zero-shot alignment of video to external audio by using a preprocessed audio envelope as the temporal condition.
+Given a waveform $y$, we compute its onset-strength envelope $s_{t}$, pool it to the number of frames, and normalize it to $[0,1]$.
+Peaks are preserved through thresholding and otherwise smoothed with a Gaussian filter:
+$\tilde{s}_{t}=\begin{cases}\hat{s}_{t},&\hat{s}_{t}\geq\tau_{audio},\\ \text{GaussianSmooth}(\hat{s}_{t};\sigma),&\text{otherwise}.\end{cases}$ (10)
+Figure 6 shows examples where video motion aligns with peaks in the audio envelope, with additional results in the appendix.
+While tested on simple cases with clear peaks, these results suggest that audio-driven temporal conditioning is a promising approach for achieving audio-video alignment without requiring training on paired datasets.
+Limitations.
+TempoControl has two main limitations.
+First, it increases inference time to approximately 460 seconds per video, compared to 170 seconds for vanilla generation.
+Nonetheless, it remains an inference-only approach that requires no additional supervision, thereby avoiding the degradation in robustness and diversity often seen in supervised training.
+Second, our conditioning can introduce mild perceptual shifts in target attributes (for example, altering the color of a cat; see Figure 1).
+While consistency is often maintained, our current objectives do not fully guarantee it, making this an interesting direction for future work.
+Temporal control in text-to-video models.
+We observe that adding explicit timing cues to current text-to-video models does not improve temporal accuracy and often degrades visual quality.
+For example, in Wan 2.1 (one-object setting), temporal accuracy remains at 64% while Imaging Quality drops from 56% to 53%.
+This highlights a broader limitation: existing models lack reliable temporal grounding from text alone.
+For fairness, we keep the same prompt format for all baselines, though omitting explicit timing slightly improves our Imaging Quality (56% vs.
+53%) without affecting Temporal Accuracy (details in the appendix).
 5.4 Conclusions
@@
 We hope our approach will inspire new solutions to tasks such as audio-video alignment, as discussed in this work.
-Supplementary Material: TempoControl – Temporal Attention Guidance for Text-to-Video Models
-This supplementary material provides detailed information supporting our main paper.
-It includes algorithmic implementation, experimental design, application-specific configurations, extended quantitative and qualitative results, and a human evaluation setup.
-Videos and code are provided as part of the supplementary materials package.
-Specifically, we include:
+Supplementary Material for TempoControl: Temporal Attention Guidance for Text-to-Video Models
+This supplementary material includes algorithmic pseudocode, ablation studies, experimental design details, application-specific configurations, extended quantitative and qualitative results, and the human-evaluation setup.
+Videos and code are also provided in the supplementary package.
+The sections of the supplementary material are as follows:
   1. 1.
-    Full algorithm and pseudocode (Section .6).
+    Algorithm pseudocode (Section .5).
   2. 2.
@@
     Additional qualitative results with sampled frames (Section .12).
-.6 Algorithm Pseudo-code
+.5 Algorithm Pseudo-code
 Algorithm 1 outlines our inference-time optimization procedure.
@@
 At each step $t$, we compute cross-attention maps from the video-text model $\operatorname{DiT}$ and derive the scalar temporal attention $a_{i}^{t}$ for each token $p_{i}$.
-The optimization loss $\mathcal{L}$ combines four components: correlation alignment $\mathcal{L}_{\text{c}}^{t}$, energy modulation $\mathcal{L}_{\text{enr}}^{t}$, entropy regularization $\mathcal{L}_{\text{ent}}^{t}$, and spatial consistency $\mathcal{L}_{\text{s}}^{t}$.
+The optimization loss $\mathcal{L}$ combines three components: correlation alignment $\mathcal{L}_{\text{c}}^{t}$, magnitude modulation $\mathcal{L}_{\text{enr}}^{t}$, entropy regularization and $\mathcal{L}_{\text{ent}}^{t}$.
 The latent code is updated via gradient descent for up to $l$ steps per denoising timestep, or until the Pearson correlation exceeds a threshold $\tau_{\text{corr}}$.
@@
 11:     $\mathcal{L}_{\ominus}^{t}\leftarrow\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}\leq\tau]\cdot a_{i,j}^{t}$
-12:     $\mathcal{L}_{\text{enr}}^{t}\leftarrow\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}$
+12:     $\mathcal{L}_{\text{mag}}^{t}\leftarrow\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}$
 13:     $\mathcal{L}_{\text{ent}}^{t}\leftarrow\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}>\tau]\cdot\mathcal{H}(\bar{A}^{t}_{j,i})$
-14:     $\mathcal{L}_{\text{s}}^{t}\leftarrow\sum_{j=1}^{T^{\prime}}\|\bar{A}^{T}_{j,i}-\bar{A}^{t}_{j,i}\|_{2}^{2}$
-15:     $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}_{\text{c}}^{t}+\lambda_{1}\mathcal{L}_{\text{enr}}^{t}+\lambda_{2}\mathcal{L}_{\text{ent}}^{t}+\lambda_{3}\mathcal{L}_{\text{s}}^{t}$
-16:   end for
-17:   $z_{t}^{\prime}\leftarrow z_{t}-\alpha_{t}\nabla_{z_{t}}(\mathcal{L}/N)$
-18:   if All $\mathcal{L}_{\text{corr}}^{t}\leq\tau_{\text{corr}}$ then
-19:     break
-20:   end if
-21:   $z_{t}\leftarrow z_{t}^{\prime}$
-22:  end for
-23:  $z_{t-1}\leftarrow\operatorname{DiT}(z_{t}^{\prime},\mathcal{P},t)$
-24:end for
-25:Continue denoising using the standard denoising diffusion process for $t=T{-}k,\dots,1$
-26:return $z_{0}$
-.7 Experimental Setup
-Hyperparameters were tuned early using a small validation set specific to each task.
-While large-scale ablations were limited by computational constraints, we found the method to be robust across scenarios with only minor adjustments.
-All experiments are conducted using Wan 2.1 as the backbone. We apply a classifier-free guidance scale of 6 and a sample shift of 3 during generation.
-For single-object scenes, we use $\lambda_{1}{=}1$, $\lambda_{2}{=}10$, $\lambda_{3}{=}1$, a learning rate of $0.0005$, a Pearson stopping threshold of $\tau_{\text{corr}}{=}0.9$, and apply optimization for the first l=10l{=}10 denoising steps with up to k=10k{=}10 gradient updates per step.
-In two-object scenes, which require more aggressive guidance, we raise the learning rate to $0.001$ for improved convergence.
-For motion-centric prompts, we reduce the temporal alignment weight to $\lambda_{1}{=}0.3$ and relax the Pearson threshold to $\tau_{\text{corr}}{=}0.85$ to allow more flexible optimization.
-.8 Applications Setup
+14:     $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}_{\text{c}}^{t}+\lambda_{1}\mathcal{L}_{\text{enr}}^{t}+\lambda_{2}\mathcal{L}_{\text{ent}}^{t}$
+15:   end for
+16:   $z_{t}^{\prime}\leftarrow z_{t}-\alpha_{t}\nabla_{z_{t}}(\mathcal{L}/N)$
+17:   if All $\mathcal{L}_{\text{corr}}^{t}\leq\tau_{\text{corr}}$ then
+18:     break
+19:   end if
+20:   $z_{t}\leftarrow z_{t}^{\prime}$
+21:  end for
+22:  $z_{t-1}\leftarrow\operatorname{DiT}(z_{t}^{\prime},\mathcal{P},t)$
+23:end for
+24:Continue denoising using the standard denoising diffusion process for $t=T{-}k,\dots,1$
+25:return $z_{0}$
+.6 Experimental Setup
+Hyperparameters were tuned early on using a small validation split tailored to each task.
+Although large-scale ablations were constrained by computational cost, the method proved robust across settings, requiring only minor adjustments.
+All experiments use Wan 2.1 as the backbone model. We apply a classifier-free guidance scale of 6 and a sample shift of 3 during generation.
+For single-object scenes, we use $\lambda_{1}{=}0.3$, $\lambda_{2}{=}10$, a learning rate of $5{\times}10^{-4}$, a Pearson stopping threshold of $\tau_{\text{corr}}{=}0.9$, and perform optimization over the first l=10l{=}10 denoising steps with up to k=10k{=}10 gradient updates per step.
+For two-object scenes, which demand stronger guidance, we increase the learning rate to $10^{-3}$ to improve convergence.
+For motion-centric prompts, we relax the Pearson threshold to $\tau_{\text{corr}}{=}0.85$ to enable more flexible optimization.
+.7 Applications Setup
 Single-Object Temporal Control.
-To assess temporal grounding for individual objects, we construct a dataset using the 80 object categories detectable by the YOLO detector.
-These span common classes, such as animals and vehicles, as well as specific items, including handbags and sports equipment.
-Videos are generated from prompts of the form: “An empty scene.
-Suddenly, $\langle o_{1}\rangle$ appears out of nowhere, drawing all attention.” Object classes are uniformly distributed across different time intervals.
+To evaluate temporal grounding for individual objects, we construct a dataset based on the 80 object categories detectable by the YOLO detector.
+These categories cover common entities such as animals, vehicles, household items, and everyday objects.
+The complete class list is provided below:
+person, bicycle, car, motorcycle, airplane, bus, train, truck, boat, traffic light, fire hydrant, stop sign, parking meter, bench, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, backpack, umbrella, handbag, tie, suitcase, frisbee, skis, snowboard, sports ball, kite, baseball bat, baseball glove, skateboard, surfboard, tennis racket, bottle, wine glass, cup, fork, knife, spoon, bowl, banana, apple, sandwich, orange, broccoli, carrot, hot dog, pizza, donut, cake, chair, couch, potted plant, bed, dining table, toilet, tv, laptop, mouse, remote, keyboard, cell phone, microwave, oven, toaster, sink, refrigerator, book, clock, vase, scissors, teddy bear, hair drier, toothbrush.
+Videos are generated from prompts of the form ‘An empty scene.
+Suddenly, $\langle o_{1}\rangle$ appears out of nowhere, drawing all attention.’ Object classes are uniformly sampled across different temporal intervals.
 Two-Object Temporal Control.
-We evaluate TempoControl on 82 object pairs from the VBench protocol.
-Prompts take the form: “The video begins with a serene view centered on $\langle o_{1}\rangle$, with no sign of $\langle o_{2}\rangle$.
-In the second half, $\langle o_{2}\rangle$ unexpectedly appears, altering the dynamic of the scene.” A static control signal (all ones) is used for $\langle o_{1}\rangle$.
-We sample 20 evenly spaced frames and require $\langle o_{1}\rangle$ to always be present.
+We evaluate TempoControl on 82 prompts derived from the VBench protocol, each consisting of a pair of objects. The complete numbered list of pairs is provided below:
+  1. 1.
+    a bird and a cat
+  2. 2.
+    a cat and a dog
+  3. 3.
+    a dog and a horse
+  4. 4.
+    a horse and a sheep
+  5. 5.
+    a sheep and a cow
+  6. 6.
+    a cow and an elephant
+  7. 7.
+    an elephant and a bear
+  8. 8.
+    a bear and a zebra
+  9. 9.
+    a zebra and a giraffe
+  10. 10.
+    a giraffe and a bird
+  11. 11.
+    a chair and a couch
+  12. 12.
+    a couch and a potted plant
+  13. 13.
+    a potted plant and a tv
+  14. 14.
+    a tv and a laptop
+  15. 15.
+    a laptop and a remote
+  16. 16.
+    a remote and a keyboard
+  17. 17.
+    a keyboard and a cell phone
+  18. 18.
+    a cell phone and a book
+  19. 19.
+    a book and a clock
+  20. 20.
+    a clock and a backpack
+  21. 21.
+    a backpack and an umbrella
+  22. 22.
+    an umbrella and a handbag
+  23. 23.
+    a handbag and a tie
+  24. 24.
+    a tie and a suitcase
+  25. 25.
+    a suitcase and a vase
+  26. 26.
+    a vase and scissors
+  27. 27.
+    scissors and a teddy bear
+  28. 28.
+    a teddy bear and a frisbee
+  29. 29.
+    a frisbee and skis
+  30. 30.
+    skis and a snowboard
+  31. 31.
+    a snowboard and a sports ball
+  32. 32.
+    a sports ball and a kite
+  33. 33.
+    a kite and a baseball bat
+  34. 34.
+    a baseball bat and a baseball glove
+  35. 35.
+    a baseball glove and a skateboard
+  36. 36.
+    a skateboard and a surfboard
+  37. 37.
+    a surfboard and a tennis racket
+  38. 38.
+    a tennis racket and a bottle
+  39. 39.
+    a bottle and a chair
+  40. 40.
+    an airplane and a train
+  41. 41.
+    a train and a boat
+  42. 42.
+    a boat and an airplane
+  43. 43.
+    a bicycle and a car
+  44. 44.
+    a car and a motorcycle
+  45. 45.
+    a motorcycle and a bus
+  46. 46.
+    a bus and a traffic light
+  47. 47.
+    a traffic light and a fire hydrant
+  48. 48.
+    a fire hydrant and a stop sign
+  49. 49.
+    a stop sign and a parking meter
+  50. 50.
+    a parking meter and a truck
+  51. 51.
+    a truck and a bicycle
+  52. 52.
+    a toilet and a hair drier
+  53. 53.
+    a hair drier and a toothbrush
+  54. 54.
+    a toothbrush and a sink
+  55. 55.
+    a sink and a toilet
+  56. 56.
+    a wine glass and a chair
+  57. 57.
+    a cup and a couch
+  58. 58.
+    a fork and a potted plant
+  59. 59.
+    a knife and a tv
+  60. 60.
+    a spoon and a laptop
+  61. 61.
+    a bowl and a remote
+  62. 62.
+    a banana and a keyboard
+  63. 63.
+    an apple and a cell phone
+  64. 64.
+    a sandwich and a book
+  65. 65.
+    an orange and a clock
+  66. 66.
+    broccoli and a backpack
+  67. 67.
+    a carrot and an umbrella
+  68. 68.
+    a hot dog and a handbag
+  69. 69.
+    a pizza and a tie
+  70. 70.
+    a donut and a suitcase
+  71. 71.
+    a cake and a vase
+  72. 72.
+    an oven and scissors
+  73. 73.
+    a toaster and a teddy bear
+  74. 74.
+    a microwave and a frisbee
+  75. 75.
+    a refrigerator and skis
+  76. 76.
+    a bicycle and an airplane
+  77. 77.
+    a car and a train
+  78. 78.
+    a motorcycle and a boat
+  79. 79.
+    a person and a toilet
+  80. 80.
+    a person and a hair drier
+  81. 81.
+    a person and a toothbrush
+  82. 82.
+    a person and a sink
+Prompts follow the structure: ‘The video begins with a serene view centered on $\langle o_{1}\rangle$, with no sign of $\langle o_{2}\rangle$.
+In the second half, $\langle o_{2}\rangle$ unexpectedly appears, altering the dynamic of the scene.’
+A static control signal (all ones) is used for $\langle o_{1}\rangle$.
+We sample 20 evenly spaced frames and require $\langle o_{1}\rangle$ to be present throughout.
 A frame is counted as successful if both objects are detected when $\langle o_{2}\rangle$ is active, and only $\langle o_{1}\rangle$ is detected when $\langle o_{2}\rangle$ is inactive.
 Motion Temporal Control.
-To test fine-grained temporal modulation of actions, we generate 100 videos using prompts such as: “A video of $\langle s\rangle$ $\langle v\rangle$ $\langle a\rangle$ with a strong movement at the $\langle t\rangle$ second,” where $\langle s\rangle$, $\langle v\rangle$, $\langle a\rangle$, and $\langle t\rangle$ denote subject, verb, adverb, and time reference, respectively.
-Example: “A child jumping joyfully at the third second.” The verb is controlled via a binary temporal mask (1 only at $\langle t\rangle$), while the subject is always present.
-Multi-Object Accuracy.
-We benchmark TempoControl on the VBench Multiple Objects benchmark, which evaluates the generation of all the mentioned objects.
-Each video is sampled at 16 evenly spaced frames.
-A frame is successful if both target objects are detected.
-The final score is the average per-frame success rate across all videos.
-We use all-one control signals for both objects and optimize only the first 4 denoising steps, skipping the Pearson-based stopping criterion.
-.9 Human Evaluation Protocol
+To evaluate fine-grained temporal modulation of actions, we generate 100 videos using prompts of the form: ‘A video of $\langle s\rangle$ $\langle v\rangle$ $\langle a\rangle$ with a strong movement at the $\langle t\rangle$ second,’ where $\langle s\rangle$, $\langle v\rangle$, $\langle a\rangle$, and $\langle t\rangle$ denote the subject, verb, adverb, and time reference, respectively.
+The full list of subject–verb–adverb triplets is given below:
+  1. 1.
+    a cat pouncing silently
+  2. 2.
+    a dancer twirling elegantly
+  3. 3.
+    a bird flapping frantically
+  4. 4.
+    a lion roaring loudly
+  5. 5.
+    a wave crashing violently
+  6. 6.
+    a ballerina leaping lightly
+  7. 7.
+    a helicopter spinning noisily
+  8. 8.
+    a snake slithering smoothly
+  9. 9.
+    a bear growling deeply
+  10. 10.
+    a man running fast
+  11. 11.
+    a woman laughing brightly
+  12. 12.
+    a child crawling slowly
+  13. 13.
+    a monkey swinging playfully
+  14. 14.
+    a firework exploding suddenly
+  15. 15.
+    a wolf howling hauntingly
+  16. 16.
+    a dolphin diving swiftly
+  17. 17.
+    a plane ascending steadily
+  18. 18.
+    a kangaroo hopping energetically
+  19. 19.
+    a fox darting quickly
+  20. 20.
+    a shadow moving mysteriously
+  21. 21.
+    a squirrel scampering nervously
+  22. 22.
+    a truck crashing loudly
+  23. 23.
+    a child screaming unexpectedly
+  24. 24.
+    a penguin waddling adorably
+  25. 25.
+    a frog croaking rhythmically
+  26. 26.
+    a leaf drifting slowly
+  27. 27.
+    a shark circling menacingly
+  28. 28.
+    a comet streaking brightly
+  29. 29.
+    a dancer collapsing dramatically
+  30. 30.
+    a boxer punching fiercely
+  31. 31.
+    a tree swaying gently
+  32. 32.
+    a man lifting weights powerfully
+  33. 33.
+    a windmill rotating steadily
+  34. 34.
+    a girl skipping cheerfully
+  35. 35.
+    a car drifting dangerously
+  36. 36.
+    a train accelerating fast
+  37. 37.
+    a snake striking quickly
+  38. 38.
+    a fire burning intensely
+  39. 39.
+    a glacier cracking slowly
+  40. 40.
+    a bee buzzing constantly
+  41. 41.
+    a deer sprinting fearfully
+  42. 42.
+    a volcano erupting violently
+  43. 43.
+    a runner collapsing from exhaustion
+  44. 44.
+    a gymnast flipping smoothly
+  45. 45.
+    a rocket launching thunderously
+  46. 46.
+    a drummer hitting rapidly
+  47. 47.
+    a magician vanishing mysteriously
+  48. 48.
+    a spider crawling delicately
+  49. 49.
+    a tiger pacing restlessly
+  50. 50.
+    a surfer balancing skillfully
+  51. 51.
+    a leopard growling softly
+  52. 52.
+    a swimmer diving gracefully
+  53. 53.
+    a baby clapping happily
+  54. 54.
+    a mime gesturing expressively
+  55. 55.
+    a crow cawing sharply
+  56. 56.
+    a goat headbutting suddenly
+  57. 57.
+    a girl blowing bubbles gently
+  58. 58.
+    a chef chopping rapidly
+  59. 59.
+    a horse shaking its mane proudly
+  60. 60.
+    a robot malfunctioning erratically
+  61. 61.
+    a meteor falling fast
+  62. 62.
+    a snake coiling tightly
+  63. 63.
+    a rabbit thumping nervously
+  64. 64.
+    a man collapsing dramatically
+  65. 65.
+    a car braking suddenly
+  66. 66.
+    a hawk diving precisely
+  67. 67.
+    a wolf snarling viciously
+  68. 68.
+    a plane landing smoothly
+  69. 69.
+    a firetruck speeding urgently
+  70. 70.
+    a chimpanzee clapping playfully
+  71. 71.
+    a dancer stomping rhythmically
+  72. 72.
+    a snail inching slowly
+  73. 73.
+    a lioness crouching quietly
+  74. 74.
+    a storm cloud rolling ominously
+  75. 75.
+    a bee landing precisely
+  76. 76.
+    a magician pulling a rabbit suddenly
+  77. 77.
+    a glacier melting steadily
+  78. 78.
+    a jellyfish floating gracefully
+  79. 79.
+    a tree falling loudly
+  80. 80.
+    a drummer drumming fiercely
+  81. 81.
+    a torch blazing brightly
+  82. 82.
+    a woman spinning dramatically
+  83. 83.
+    a child running barefoot
+  84. 84.
+    a fox sniffing cautiously
+  85. 85.
+    a man throwing a ball forcefully
+  86. 86.
+    a bear shaking off water heavily
+  87. 87.
+    a lion walking regally
+  88. 88.
+    a horse rearing suddenly
+  89. 89.
+    a falcon gliding silently
+  90. 90.
+    an elephant raising its trunk highly
+  91. 91.
+    a dog barking forcefully
+  92. 92.
+    a man sneezing powerfully
+  93. 93.
+    a child jumping joyfully
+  94. 94.
+    a tiger leaping fiercely
+  95. 95.
+    a woman spinning gracefully
+  96. 96.
+    a robot marching stiffly
+  97. 97.
+    a horse galloping wildly
+  98. 98.
+    an eagle soaring majestically
+  99. 99.
+    a flame flickering rapidly
+  100. 100.
+    a skateboarder flipping expertly
+The verb is modulated using a binary temporal mask (set to 1 only at time $\langle t\rangle$), while the subject remains active throughout the video.
+.8 Human Evaluation Protocol
 To assess temporal alignment and visual quality, we conducted a human preference study using side-by-side video comparisons.
-For each prompt, two videos were presented, one generated by our method and one by the text baseline, randomly assigned as Video A or B to avoid bias.
-Annotators were asked to read the prompt and answer two questions: (1) Which video is more temporally accurate (i.e., does the object appear at the correct time and remain visible)?
-and (2) Which video looks more visually appealing?
-A “Same (Use sparingly!)” option was available if no difference was perceived.
-Example screenshots of the evaluation interface are shown in Figures 6 and 7.
+All annotators were graduate students conducting research in computer vision or closely related areas.
+For each prompt, two videos were shown: one generated by our method and one by the text-only baseline, with their order (Video A or Video B) randomized to avoid bias.
+Annotators were instructed to read the prompt and answer two questions: (1) Which video is more temporally accurate (i.e., does the object appear at the correct time and remain visible)?
+(2) Which video is more visually appealing?
+A “Same (use sparingly)” option was available when no meaningful difference was perceived.
+After the experiments, several annotators noted that judging temporal control could be challenging when the difference in timing was only 1–2 seconds.
+They reported that knowing the exact expected trigger time from the control mask would have helped them assess temporal accuracy more reliably.
+While we chose not to repeat the study, we include this observation as guidance for future temporal evaluation protocols.
+Example screenshots of the evaluation interface are shown in Figures 7 and 8.
 Figure 7: Evaluation interface showing the prompt and two videos (A and B) for comparison.
 Figure 8: Second example of the human evaluation interface.
-.10 Multi-Object Evaluation
-We report extended results on multi-object text-to-video generation using the VC-Bench benchmark.
-In this setting, prompts reference multiple entities with temporal conditions (e.g., “A dog in the beginning, then a cat”).
-We compare our method against a text baseline.
-Table 4 summarizes results.
-Our method achieves substantial gains in object-centric metrics.
-Specifically, we improve multi-object accuracy from 74.1% to 76.4% (GriT) and from 61.5% to 65.7% (YOLO), demonstrating improved temporal grounding and visibility of both entities.
-We also observe improvements across all consistency metrics, including subject consistency (from 97.2% to 97.8%) and background consistency (from 97.6% to 98.1%).
-The dynamic degree shows a slight reduction (from 30.5% to 18.8%), which reflects a known tradeoff: our method promotes temporal consistency, often leading to more stable object appearance throughout the video.
-As a result, motion may be reduced.
-However, we note a limitation of this metric: VC-Bench implicitly favors videos where objects are present in all frames.
-A more suitable benchmark might instead reward appearances rather than requiring full-video presence.
+.9 Multi-Object Evaluation
+We report extended results on multi-object text-to-video generation using the VBench benchmark.
+In this setting, prompts reference multiple entities (e.g., ‘A dog and a cat’), and the evaluation measures whether all specified objects appear consistently throughout the video.
+Each video is sampled at 16 evenly spaced frames, and a frame is counted as successful if both target objects are detected; the final score is the average per-frame success rate.
+In the context of our method, this setup corresponds to applying a constant mask of one for both objects across all frames.
+For completeness, we briefly summarize the VBench metrics reported in Table 5:
+  • Multiple-Objects Accuracy (GriT / YOLO).
+Measures whether all referenced objects appear correctly in each sampled frame, using open-vocabulary detectors (GriT) or standard detectors (YOLO).
+Higher values indicate more reliable object visibility and grounding.
+  • Subject Consistency.
+Quantifies how consistently the appearance of the main subject is preserved across frames.
+It is computed using feature similarity both between consecutive frames and relative to the first frame of the video; higher values indicate a more stable and coherent subject identity.
+  • Background Consistency. Measures the stability of the background across the video. High scores indicate coherent scene structure without abrupt changes.
+  • Dynamic Degree.
+Captures the amount of motion or temporal change in the video.
+The metric is binary at the video level: a video is marked as dynamic if its average optical flow surpasses a predefined threshold.
+Higher values indicate more motion, while lower values reflect greater temporal stability.
+Table 5 summarizes the results.
+Our method achieves substantial gains in object-centric metrics: multi-object accuracy improves from 74.1% to 76.4% (GriT) and from 61.5% to 65.7% (YOLO), demonstrating better grounding and more reliable visibility of both entities.
+We also observe improvements across consistency metrics, including subject consistency (from 97.2% to 97.8%) and background consistency (from 97.6% to 98.1%).
+The Dynamic Degree decreases from 30.5% to 18.8%, reflecting a trade-off: our method promotes temporal stability, often leading to more consistent object appearance across the video, at the cost of reduced motion.
+This highlights a limitation of this evaluation protocol: VBench counts an object as correct only when it is present in every sampled frame; hence, the benchmark implicitly favors unnatural videos in which objects persist throughout the sequence.
 Table 5: Full results for Multiple object benchmark results. Best results per column are in bold.
 Method Multiple Object (GriT) Multiple Object (YOLO) Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality Text 74.13% 61.54% 97.22% 97.56% 99.18% 30.49% 62.84% 70.25% Ours 76.37% 65.73% 97.81% 98.10% 99.40% 18.78% 62.52% 70.21%
-.11 Ablations
-Table 5 presents an ablation study in the two-object setting, evaluating the impact of removing the energy term, entropy term, and spatial penalty.
-Omitting the energy term results in a modest decrease in temporal accuracy (from 55% to 52.56%).
-In contrast, removing the entropy term has a more significant effect: temporal accuracy falls to 48.96%, and the temporal presence accuracy drops sharply from 43.54% to 34.76%.
-Subject Consistency, Aesthetic Quality, and Imaging Quality also decline.
-Without entropy regularization, attention becomes overly diffuse, often blending the two objects.
-Qualitative analysis shows that without the entropy term, the model frequently replaces the static object with the dynamic one (see Section .8 for details on the static and temporally conditioned objects, and Figure 8 for qualitative results).
-Table 5: Ablation results for the two-object setting.
-Method Temporal Accuracy T.
-Absence Accuracy T.
-Presence Accuracy Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality Text baseline 37.50% 45.85% 29.15% 92.06% 95.32% 98.92% 62.20% 53.26% 68.56% Ours 55.00% 66.46% 43.54% 92.21% 95.43% 99.19% 41.46% 56.55% 70.66% No energy 52.56% 58.90% 46.22% 92.31% 95.45% 99.34% 46.34% 56.12% 71.15% No entropy 48.96% 63.17% 34.76% 89.99% 94.39% 98.92% 53.66% 54.66% 68.95% No spatial 52.01% 60.85% 43.17% 92.76% 95.67% 99.23% 48.78% 56.75% 70.25%
-Figure 8: Ablation study. Left: without entropy regularization, the motorcycle disappears. Right: the bear is incorrectly replaced by a zebra.
-.12 Impact of Temporal Phrasing
-We analyze the effect of using explicit temporal phrasing (e.g., “in the third second,” “in the second part of the video”) in the prompt for the one-object setup.
-Table 6 compares Wan 2.1 and our method, each evaluated with and without such phrasing.
-For Wan 2.1, omitting temporal cues slightly improves temporal accuracy (63.94% to 65.18%) and yields higher aesthetic (45.98% to 48.78%) and imaging quality (53.76% to 59.99%), suggesting that the model struggles to interpret or apply explicit timing effectively.
-Our method benefits even more: removing temporal phrases results in the highest temporal accuracy (81.12%) along with gains in aesthetic (48.16%) and imaging quality (56.32%).
-We hypothesize that Wan 2.1 is not trained to handle precise temporal expressions, and current attention mechanisms may not effectively capture their semantics.
-As a result, such phrasing introduces ambiguity.
-In contrast, concise prompts without explicit timing allow our method to align more reliably with the intended temporal structure.
-For fairness, all main paper results use prompts with temporal phrasing to match the baseline.
-However, these findings suggest that our method achieves even stronger performance when using simpler prompts without explicit timing constraints.
+.10 Ablations
+We conduct an ablation study to analyze the influence of the magnitude weight $\lambda_{1}$ and the entropy weight $\lambda_{2}$ on temporal control and visual quality.
+The results reveal clear and complementary effects of the two loss components.
+Increasing the magnitude term $\lambda_{1}$ reduces Imaging Quality while improving T.
+Absence Accuracy.
+This suggests that stronger magnitude guidance more strictly enforces the temporal mask, but may also reduce overall visual fidelity.
+In contrast, increasing the entropy term $\lambda_{2}$ improves Imaging Quality and raises T.
+Presence Accuracy, but it decreases T.
+Absence Accuracy, indicating that stronger entropy regularization encourages more persistent object appearance.
+Overall, our chosen hyperparameters ($\lambda_{1}=0.3$, $\lambda_{2}=10$) provide a balanced tradeoff between these competing effects, achieving strong Temporal Accuracy while maintaining high Imaging Quality.
+Table 6: Ablation results for the one object setting.
+All variants use the same temporal Pearson term; we ablate the magnitude weight $\lambda_{1}$ and the entropy weight $\lambda_{2}$.
+Best scores per column are in bold.
+(a) Effect of magnitude weight $\lambda_{1}$ with fixed entropy weight $\lambda_{2}=10$.
+Method
+Temporal
+Accuracy
+T. Absence
+Accuracy
+T. Presence
+Accuracy
+Imaging
+Quality
+Text baseline 63.94% 67.38% 60.50% 53.76%
+Ours ($\lambda_{1}{=}0.3,\lambda_{2}{=}10$) 82.50% 83.25% 81.75% 56.51%
+$\lambda_{1}{=}0.1,\lambda_{2}{=}10$ 79.50% 80.62% 78.38% 56.37%
+$\lambda_{1}{=}0.65,\lambda_{2}{=}10$ 81.06% 86.50% 75.62% 53.94%
+$\lambda_{1}{=}1.0,\lambda_{2}{=}10$ 83.12% 89.00% 77.25% 53.74%
+(b) Effect of entropy weight $\lambda_{2}$ with fixed magnitude weight $\lambda_{1}=0.3$.
+Method
+Temporal
+Accuracy
+T. Absence
+Accuracy
+T. Presence
+Accuracy
+Imaging
+Quality
+Text baseline 63.94% 67.38% 60.50% 53.76%
+Ours ($\lambda_{1}{=}0.3,\lambda_{2}{=}10$) 82.50% 83.25% 81.75% 56.51%
+$\lambda_{1}{=}0.3,\lambda_{2}{=}1$ 80.38% 91.25% 69.50% 52.01%
+$\lambda_{1}{=}0.3,\lambda_{2}{=}5$ 81.88% 88.50% 75.25% 54.04%
+$\lambda_{1}{=}0.3,\lambda_{2}{=}15$ 81.56% 81.62% 81.50% 56.31%
+Figure 9 illustrates ablation results for the entropy term, revealing the limitation of using the Pearson correlation term alone without entropy regularization.
+While Pearson-based optimization successfully forces the object to appear at the desired time, it often alters the object’s semantics, leading to noticeable inconsistencies.
+For example, in the “Cake” row, the Pearson-only variant transforms the cake into a noticeably different shape and texture; in the “Sandwich” row, the sandwich becomes distorted across frames; and in the “Book” row, the open book becomes a closed stack with altered proportions.
+Similar semantic drift is observed in the “Keyboard,” “Oven,” and “Stop sign” examples.
+These results highlight that Pearson correlation alone can enforce temporal timing but does not sufficiently constrain the appearance space, resulting in unwanted semantic changes.
+The entropy regularization helps mitigate those artifacts.
+Figure 9: Comparison between Wan2.1 and Pearson-only optimization for single-object temporal control.
+Pearson correlation alone often satisfies the temporal constraint but alters object semantics, producing distorted or inconsistent appearances (e.g., the cake becoming misshaped, the sandwich warping, the book changing structure, or the keyboard turning into a laptop).
+Adding entropy regularization prevents these failures by stabilizing the object’s identity across time.
+.11 Limitations of Explicit Temporal Cues in State-of-the-Art Text-to-Video Models
+We analyze the effect of using explicit temporal phrasing (e.g., ‘in the third second’, ‘in the second part of the video’) in the prompt for the one-object setup.
+Table 7 compares Wan 2.1, Wan 2.2, and our method, each evaluated with and without such phrasing.
+For Wan 2.1, explicit temporal cues do not improve temporal grounding.
+In fact, removing them increases Temporal Accuracy (63.94% to 65.18%) and leads to a clear gain in Imaging Quality (53.76% to 59.99%), indicating that the model struggles to interpret precise timing expressions and performs better without them.
+Wan 2.2 shows a more nuanced pattern: it handles temporal phrasing better than Wan 2.1, likely due to improved data or training techniques.
+However, the drop in quality when explicit timing is introduced is still significant.
+Imaging Quality decreases sharply from 57.78% (without time) to 48.13% (with time), showing that temporal phrasing introduces confusion even for stronger models.
+Presence Accuracy also falls (79.50% to 68%), further highlighting that explicit timing signals continue to degrade performance.
+Our method remains the most robust and accurate across settings.
+Across all conditions, our method provides the most reliable and precise temporal control, regardless of prompt style.
+Interestingly, although our approach builds on Wan 2.1 as the base model, explicit temporal phrasing does increase the temporal accuracy.
+This suggests that our optimization procedure may curate the base model’s timing capabilities.
 Table 7: Effect of temporal phrasing in prompts for the one-object setup. Removing explicit timing improves video quality.
-Method Temp.
-Acc.
-T.
-Abs.
-Acc.
-T.
-Pres.
-Acc.
-Subj.
-Cons.
-Backg.
-Cons.
-Motion Smooth.
-Dyn.
-Deg.
-Aesthetic Imaging Wan with time 63.94% 67.38% 60.50% 84.66% 92.30% 99.16% 47.50% 45.98% 53.76% Wan without time 65.18% 65.12% 65.25% 85.29% 92.44% 99.22% 45.00% 48.78% 59.99% Ours with time 81.00% 88.75% 72.91% 80.60% 90.33% 99.05% 38.75% 44.37% 52.99% Ours without time 81.12% 84.75% 77.50% 81.29% 90.72% 99.06% 35.00% 48.16% 56.32%
-.13 Additional Qualitative Examples
+Method Temp. Acc. T. Abs. Acc. T. Pres. Acc. Imaging Quality
+Wan2.1 with time 63.94% 67.38% 60.50% 53.76%
+Wan2.1 without time 65.18% 65.12% 65.25% 59.99%
+Wan2.2 with time 78.02% 87.42% 68.00% 48.13%
+Wan2.2 without time 75.24% 71.25% 79.50% 57.78%
+Ours with time 82.50% 83.25% 81.75% 56.51%
+Ours without time 78.75% 82.25% 75.25% 57.14%
+.12 Additional Qualitative Examples
 We provide qualitative results across the one-object, two-object, and motion-based setups.
@@
 Videos can be viewed on the project page.
+a baseball glove and a skateboard
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a bicycle and a car
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a bicycle and an airplane
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a bird and a cat
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a boat and an airplane
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a book and a clock
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a bowl and a remote
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a bus and a traffic light
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a cake and a vase
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a car and a motorcycle
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a cat and a dog
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a cell phone and a book
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a donut and a suitcase
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Figure 10: Two Objects Text vs Ours.
 TextOurs
-apple 4sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+a giraffe and a bird
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a handbag and a tie
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a horse and a sheep
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a laptop and a remote
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a motorcycle and a boat
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a motorcycle and a bus
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a person and a toothbrush
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a pizza and a tie
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a potted plant and a tv
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a spoon and a laptop
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a tennis racket and a bottle
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a toaster and a teddy bear
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a truck and a bicycle
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+a zebra and a giraffe
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Figure 11: Two Objects Text vs Ours.
 TextOurs
-bear 3sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-bench 2sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-bus 2sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-clock 5sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-elephant 3sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-surfboard 3sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-tie 3sec
-TextOurs
-umbrella 3sec
+airplane 2nd second
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-vase 5sec
+apple 4th second
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-zebra 3sec
+backpack 3rd second
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-wine glass 4sec
+bench 2nd second
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-kite 3sec
+bicycle 2nd second
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-fire hydrant 2sec
+bowl 4th second
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-skateboard 3sec
+bus 2nd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+cell phone last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+clock last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+cow 2nd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+dining table last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+dog last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+donut 4th second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+fork 4th second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
 Figure 12: One Object Text vs Ours.
 TextOurs
-a bird and a cat
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+hair drier last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+handbag 3rd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+kite 3rd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+knife 4th second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+scissors last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+sheep 2nd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+skateboard 3rd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+snowboard 3rd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+sports ball 3rd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+suitacase 3rd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+toilet last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+truck 2nd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+umbrella 3rd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Figure 13: One Object Text vs Ours.
 TextOurs
-a boat and an airplane
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+bee buzzing 3rd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+bird flapping 2nd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+child crawling 2nd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+child jumping 1st second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+child screaming 2nd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+chimpanze clapping 1st second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+crow cawing 4th second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+falcon gliding last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+fox darting 2nd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+horse shaking its mane 4th second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+jellyfish floating last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+kangaroo hopping 2nd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+leopard growling 4th second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Figure 14: Movement Text vs Ours.
 TextOurs
-a book and a clock
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+lion roaring 1st second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+lion walking last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+magician pulling rabbit last second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+penguin waddling 1st second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+rabbit thumping 4th second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+robot marching 1st second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+rocket launching 3rd second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+surfer balancing 4th second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+swimmer diving 4th second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+wolf snarling 4th second
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Figure 15: Movement Text vs Ours.
 TextOurs
-a carrot and an umbrella
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-a chair and a couch
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-a giraffe and a bird
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-a horse and a sheep
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-a teddy bear and a frisbee
-TextOurs
-a train and a boat
+elephant ex1
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-a truck and a bicycle
+elephant ex2
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-a wine glass and a chair
+lightning ex1
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-a zebra and a giraffe
+lightning ex2
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-an umbrella and a handbag
-TextOurs
-a bear and a zebra
+rooster ex1
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-a cow and an elephant
-Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
-TextOurs
-a cat and a dog
-Figure 10: Two objects Text vs Ours.
-TextOurs
-chimpanze clapping 1sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-horse shaking its mane 4sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-lion roaring 1sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-magician pulling rabbit 5sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-man lifting weights 3sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-rabbit thumping 4second
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-snake slithering 1sec
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-TextOurs
-tiger leaping 1sec
-TextOurs
-child jumping 1sec
+rooster ex2
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOurs
-wolf howling 2sec
-Figure 11: Movement Text vs Ours.
+Figure 16: Audio-video Alignment Text vs Ours.